### v1

Abstract
In this paper, we study the Eulerian Strong Component Arc Deletion problem, where the input is a directed multigraph and the goal is to delete the minimum number of arcs to ensure every strongly connected component of the resulting digraph is Eulerian.
This problem is a natural extension of the Directed Feedback Arc Set problem and is also known to be motivated by certain scenarios arising in the study of housing markets. The complexity of the problem, when parameterized by solution size (i.e., size of the deletion set), has remained unresolved and has been highlighted in several papers. In this work, we answer this question by ruling out (subject to the usual complexity assumptions) a fixed-parameter tractable (FPT) algorithm for this parameter and conduct a broad analysis of the problem with respect to other natural parameterizations. We prove both positive and negative results. Among these, we demonstrate that the problem is also hard (W[1]-hard or even para-NP-hard) when parameterized by either treewidth or maximum degree alone. Complementing our lower bounds, we establish that the problem is in XP when parameterized by treewidth and FPT when parameterized either by both treewidth and maximum degree or by both treewidth and solution size. We show that these algorithms have near-optimal asymptotic dependence on the treewidth assuming the Exponential Time Hypothesis.
1 Introduction
In the Eulerian Strong Component Arc Deletion (ESCAD) problem, where the input is a directed graph (digraph) [^2] In this paper, the arc set of a digraph is a multiset, i.e., we allow multiarcs. Moreover, we treat multiarcs between the same ordered pairs of vertices as distinct arcs in the input representation of all digraphs. Consequently, the number of arcs in the input is upper bounded by the length of the input. We exclude loops as they play no non-trivial role in instances of this problem. and a number $k$ and the goal is to delete at most $k$ arcs to ensure every strongly connected component of the resulting digraph is Eulerian. This problem was first introduced by Cechlárová and Schlotter [2] to model problems arising in the study of housing markets and they left the existence of an FPT algorithm for ESCAD as an open question.
The ESCAD problem extends the well-studied Directed Feedback Arc Set (DFAS) problem. In DFAS, the goal is to delete the minimum number of arcs to make the digraph acyclic. The natural extension of DFAS to ESCAD introduces additional complexity as we aim not to prevent cycles, but aim to balance in-degrees and out-degrees within each strongly connected component. As a result, the balance requirement complicates the problem significantly and the ensuing algorithmic challenges have been noted in multiple papers [2, 5, 10].
Crowston et al. [3] made partial progress on the problem by showing that ESCAD is fixed-parameter tractable (FPT) on tournaments and also gave a polynomial kernelization. However, the broader question of fixed-parameter tractability of ESCAD on general digraphs has remained unresolved.
Our contributions. Our first main result rules out the existence of an FPT algorithm for ESCAD under the solution-size parameterization, subject to standard complexity-theoretic assumptions.
Theorem 1.1.
ESCAD is W$[1]$-hard parameterized by the solution size.
The above negative result explains, in some sense, the algorithmic challenges encountered in previous attempts at showing tractability and shifts the focus toward alternative parameterizations. However, even here, we show that a strong parameterization such as the vertex cover number is unlikely to lead to a tractable outcome.
Theorem 1.2.
ESCAD is W$[1]$-hard parameterized by the vertex cover number of the graph.
In fact, assuming the Exponential Time Hypothesis (ETH), we are able to obtain a stronger lower bound.
Theorem 1.3.
There is no algorithm solving ESCAD in $f(k)\cdot n^{o(k/\log k)}$ time for some function $f$, where $k$ is the vertex cover number of the graph and $n$ is the input length, unless the Exponential Time Hypothesis fails.
To add to the hardness results above, we also analyze the parameterized complexity of the problem parameterized by the maximum degree of the input digraph and show that even for constant values of the parameter, the problem remains NP-hard.
Theorem 1.4.
ESCAD is NP-hard in digraphs where each vertex has $(\mathrm{in},\mathrm{out})$ degrees in $\{(1,6),(6,1)\}$.
We complement these negative results by showing that ESCAD is FPT parameterized by the treewidth of the deoriented digraph (i.e., the underlying undirected multigraph) and solution size as well as by the treewidth and maximum degree of the input digraph. Furthermore, we give an XP algorithm parameterized by treewidth alone. All three results are obtained by a careful analysis of the same algorithm stated below.
Theorem 1.5.
An ESCAD instance $\mathcal{I}=(G,k)$ can be solved in time $2^{{\mathcal{O}}({\mathtt{tw}}^{2})}\cdot(2\alpha+1)^{2{\mathtt{tw}}}\cdot n^{{\mathcal{O}}(1)}$ where ${\mathtt{tw}}$ is the treewidth of deoriented $G$, $\Delta$ is the maximum degree of $G$, and $\alpha=\min(k,\Delta)$.
In the above statement, notice that $\alpha$ is upper bounded by the number of edges in the digraph and so, implies an XP algorithm parameterized by the treewidth with running time $2^{{\mathcal{O}}({\mathtt{tw}}^{2})}\cdot n^{{\mathcal{O}}({\mathtt{tw}})}$. Notice the running time of our algorithm asymptotically almost matches our ETH based lower bound (recall that the vertex cover number of a graph is at least the treewidth) in Theorem 1.3.
Recall that in general, multiarcs are permitted in an instance of ESCAD. This fact is crucially used in the proof of Theorem 1.2 and raises the question of adapting this reduction to simple digraphs (digraphs without multiarcs or loops) in order to obtain a similar hardness result parameterized by vertex cover number. However, we show that this is not possible by giving an FPT algorithm for the problem on simple digraphs parameterized by the vertex integrity of the input graph. Recall that a digraph has vertex integrity $k$ if there exists a set of vertices of size $q\leq k$ which when removed, results in a digraph where each weakly connected component has size at most $k-q$. Vertex integrity is a parameter lower bounding vertex cover number and has gained popularity in recent years as a way to obtain FPT algorithms for problems that are known to be W$[1]$-hard parameterized by treedepth – one example being ESCAD on simple graphs as we show in this paper (see Theorem 1.7 below).
Theorem 1.6.
ESCAD on simple digraphs is FPT parameterized by the vertex integrity of the graph.
As a consequence of this result, we infer an FPT algorithm for ESCAD on simple digraphs parameterized by the vertex cover number, highlighting the difference in the behaviour of the ESCAD problem on directed graphs that permit multiarcs versus simple digraphs. On the other hand, we show that even on simple digraphs this positive result does not extend much further to well-studied width measures such as treewidth (or even the larger parameter treedepth), by obtaining the following consequence of Theorems 1.2 and 1.3.
Theorem 1.7.
ESCAD even on simple digraphs is W$[1]$-hard parameterized by $k$ and assuming ETH, there is no algorithm solving it in $f(k)n^{(k/logk)}$ time for some function $f$, where $k$ is the size of the smallest vertex set that must be deleted from the input digraph to obtain a disjoint union of directed stars and $n$ is the input length.
Related Work. The vertex-deletion variant of ESCAD is known to be W$[1]$-hard, as shown by Göke et al. [10], who identify ESCAD as an open problem and note that gaining more insights into its complexity was a key motivation for their study. Cygan et al. [5] gave the first FPT algorithm for edge (arc) deletion to Eulerian graphs (respectively, digraphs). Here, the aim is to make the whole graph Eulerian whereas the focus in ESCAD is on each strongly connected component. Cygan et al. also explicitly highlight ESCAD as an open problem and a motivation for their work. Goyal et al. [11] later improved the algorithm of Cygan et al. by giving algorithms achieving a single-exponential dependence on $k$.
2 Preliminaries
For a digraph $G$, we denote its vertices by $V(G)$, arcs by $E(G)$, the subgraph induced by $S\subseteq V(G)$ as $G[S]$, a subgraph with subset of vertices removed as $G-S=G[V(G)\setminus S]$, and a subgraph with subset of edges $F\subseteq E(G)$ removed as $G-F=(V(G),E(G)\setminus F)$. For a vertex $v$ and digraph $G$, let $\deg^{-}_{G}(v)$ denote its in-degree, $\deg^{+}_{G}(v)$ be its out-degree, and $\deg^{+}_{G}(v)-\deg^{-}_{G}(v)$ is called its imbalance. If the imbalance of $v$ is 00 then $v$ is said to be balanced (in $G$). A digraph is called balanced if all its vertices are balanced. The maximum degree of a digraph $G$ is the maximum value of $\deg^{+}_{G}(v)+\deg^{-}_{G}(v)$ taken over every vertex $v$ in the graph.
A vertex $v$ is reachable from $u$ if there exists a directed path from $u$ to $v$ in $G$. A strongly connected component of $G$ is a maximal set of vertices where all vertices are mutually reachable. Let strong subgraph denoted $\mathrm{strong}(G)$ be the subgraph of $G$ obtained by removing all arcs that have endpoints in different strongly connected components. The ESCAD problem can now be formulated as “Is there a set $S\subseteq V(G)$ of size $|S|\leq k$ such that $\mathrm{strong}(G-S)$ is balanced?” We call an arc $e\in E(G)$ active in $G$ if $e\in E(\mathrm{strong}(G))$ and inactive in $G$ otherwise.
A graph $G$ has vertex cover $k$ if there exists a set of vertices $S\subseteq V(G)$ with bounded size $|S|\leq k$ such that $G-S$ is an independent set. A star is an undirected graph isomorphic to $K_{1}$ or $K_{1,t}$ for some $t\geq 0$ and a directed star is just a digraph whose underlying undirected graph is a star.
A tree decomposition of an undirected graph $G$ is a pair $(T,\{X_{t}\}_{t\in V(T)})$ where $T$ is a tree and $X_{t}\subseteq V(G)$ such that (i) for all edges $uv\in E(G)$ there exists a node $t\in V(T)$ such that $\{u,v\}\subseteq X_{t}$ and (ii) for all $v\in V(G)$ the subgraph induced by $\{t\in V(T):v\in X_{t}\}$ is a non-empty tree. The width of a tree decomposition is $\max_{t\in V(T)}|X_{t}|-1$. The treewidth of $G$ is the minimum width of a tree decomposition of $G$.
Let $(T,\{X_{t}\}_{t\in V(T)})$ be a tree decomposition of $G$. We refer to every node of $T$ with degree one as a leaf node except one which is chosen as the root, $r$. A tree decomposition $(T,\{X_{t}\}_{t\in V(T)})$ is a nice tree decomposition with introduce edge nodes if all of the following conditions are satisfied:
  1. 1.
    $X_{r}=\emptyset$ and $X_{\ell}=\emptyset$ for all leaf nodes $\ell$.
  2. 2.
    Every non-leaf node of $T$ is one of the following types:
    • Introduce vertex node: a node $t$ with exactly one child $t^{\prime}$ such that $X_{t}=X_{t^{\prime}}\cup v$ for some vertex $v\notin X_{t^{\prime}}$.
    • Introduce edge node: a node $t$, labeled with an edge $uv$ where $u,v\in X_{t}$ and with exactly one child $t^{\prime}$ such that $X_{t}=X_{t^{\prime}}$.
    • Forget node: a node $t$ with exactly one child $t^{\prime}$ such that $X_{t}=X_{t^{\prime}}\setminus\{v\}$ for some vertex $v\in X_{t^{\prime}}$.
    • Join node: a node $t$ with exactly two children $t_{1},t_{2}$ such that $X_{t}=X_{t_{1}}=X_{t_{2}}$.
  3. 3.
    Every edge appears on exactly one introduce edge node.
3 Our Results for ESCAD
In the following four subsections we describe three hardness results and tractability results on bounded treewidth graphs for ESCAD. In Section 3.4 we show that the problem is XP by treewidth and FPT in two cases – when parameterized by the combined parameter treewidth plus maximum degree, and when parameterized by treewidth plus solution size. The hardness results show that dropping any of these parameters leads to a case that is unlikely to be FPT. More precisely, we show that parameterized by solution size it is W$[1]$-hard (in Section 3.1) as is the case when parameterized by vertex cover number (Section 3.2), and it is para-NP-hard when when parameterized by the maximum degree (Section 3.3).
3.1 W[1]-hardness of ESCAD Parameterized by Solution Size
In this section, we show that ESCAD is W$[1]$-hard when parameterized by solution size. Our reduction is from Multicolored Clique. The input to Multicolored Clique consists of a simple undirected graph $G$, an integer $\ell$, an containing exactly one vertex from each set $V_{i},i\in[\ell]$. Multicolored Clique is known to be W$[1]$-hard when parameterized by the size of the solution $\ell$ [4]. Each set $V_{i}$ for $i\in[\ell]$ is called a color class and for a vertex $v$ in $G$, we say $v$ has color $i$ if $v\in V_{i}$. We assume without loss of generality that in the Multicolored Clique instance we reduce from, each color class $V_{i}$ forms an independent set (edges in the same color class can be removed) and moreover, for each vertex $v\in V_{i}$ and each $j\in[\ell]\setminus\{i\}$ there exists a $w\in V_{j}$ that is adjacent to $v$ (any vertex that cannot participate in a multicolored clique can be removed).
We start with descriptions of two auxiliary gadgets: the imbalance gadget and the path gadget.
Imbalance Gadget. Let $u,v$ be a pair of vertices, and $b,c$ be two positive integers. We construct a gadget $I_{u,v}$ connecting the vertex $u$ to $v$ by a path with vertices $u,w_{1},\dots,w_{b},v$ where $w_{i}$’s are $b$ new vertices (we call them intermediate vertices in this gadget); let $u=w_{0}$ and $v=w_{b+1}$. For every $i\in\{0,\dots,b\}$ the path contains $b+1+c$ forward arcs $(w_{i},w_{i+1})$ and $b+1$ backward arcs $(w_{i+1},w_{i})$, see Figure 1(a) for an illustration. Observe that with respect to the gadget $I_{u,v}$, the vertices $u$ and $v$ have imbalances $c$ and $-c$, respectively, whereas other vertices in the gadget have imbalance zero. We refer to this gadget $I_{u,v}$ as a $(b,c)$-imbalance gadget.
(a) $c$-imbalance gadget
(b) $c$-path gadget
Figure 1: Black ellipses are vertices, thick edges represent $(b+1)$ copies of the edges; $(b+2)$ is the number of vertices in the gadgets.
Path Gadget. Let $u,v$ be a pair of vertices, and $b,c$ be two positive integers. We construct a gadget $P_{u,v}$ connecting the vertex $u$ to $v$ by a path with vertices $u,w_{1},\dots,w_{b},v$ where $w_{i}$’s are $b$ new intermediate vertices; let $u=w_{0}$ and $v=w_{b+1}$. For every $i\in\{0,\dots,b\}$ the path contains $c$ forward arcs $(w_{i},w_{i+1})$. See Figure 1(b) for an illustration. Notice that, unlike the imbalance gadget, we do not add backward arcs. Observe that with respect to the gadget $P_{u,v}$, the vertices $u$ and $v$ has imbalances $c$ and $-c$, respectively, whereas the other vertices in the gadget have imbalance zero. We refer to this gadget $P_{u,v}$ as a $(b,c)$-path gadget.
We use the following properties of the gadgets $I_{uv}$ and $P_{uv}$ to reason about the correctness of our construction.
Lemma 3.1.
Let $(G,b)$ be a yes-instance of ESCAD and $S$ be a solution. Assume that for a pair of vertices $u,v$ in $G$, there is a $(b,c)$-imbalance gadget $I_{uv}$ present in $G$ (i.e., $I_{uv}$ is an induced subgraph of $G$). If $S$ is an inclusionwise minimal solution then $S$ contains no arc of $I_{uv}$.
Proof.
In the subgraph $I_{uv}$, there are $b+c+1$ arc-disjoint paths from $u$ to $v$. Hence, all the vertices in $I_{uv}$ must be contained in the same strongly connected component in $G-S$. For each $i\in[b+1]$, let $F_{i}$ be all the arcs between $w_{i-1}$ and $w_{i}$ in either direction. Assume that we have a solution $S$ that uses the minimum number of arcs from imbalance gadgets. If $S$ contains no arc of $I_{uv}$, then we are done. Otherwise, since $|S|\leq b$ and $S\cap A(I_{uv})\neq\emptyset$, there is an $i\in\{0,1,\ldots,b+1\}$ such that either $S\cap F_{i}\neq\emptyset$, $S\cap F_{i+1}=\emptyset$ or $S\cap F_{i}=\emptyset$ , $S\cap F_{i+1}\neq\emptyset$ holds. Assume that $S\cap F_{i}\neq\emptyset$, $S\cap F_{i+1}=\emptyset$. The argument in the other case is analogous. Now, consider the vertex $w_{i}$. To ensure that the imbalance of $w_{i}$ is zero in $\mathrm{strong}(G-S)$, the solution $S$ must contain the same number of out-arcs and in-arcs of $w_{i}$ from the set $F_{i}$ (due the fact that all the vertices in $I_{uv}$ must be contained in the same strongly connected component in $G-S$). Now, consider the set $S^{\prime}=S\setminus(S\cap F_{i})$. As the vertices $w_{i-1}$ and $w_{i}$ are in the same strongly connected component in $G-S$, and $S^{\prime}$ is a subset of $S$, they must also be in the same strongly connected component in $G-S^{\prime}$. We know that all the arcs in $F_{i}$ have endpoints in $w_{i-1}$ and $w_{i}$. Now, we show that in $G-S^{\prime}$, the vertices $w_{i-1}$ and $w_{i}$ remain balanced. But this is true as $S$ contains the same number of out-arcs and in-arcs of $w_{i}$ from $F_{i}$ and the other end points of those arcs must belong to $w_{i-1}$ by definition of $F_{i}$. So, $S^{\prime}$ is also a solution but uses fewer edges from $I_{uv}$ compared to $S$, contradicting the choice of $S$. ∎
Lemma 3.2.
Let $(G,b)$ be a yes-instance of ESCAD and $S$ be an inclusionwise minimal solution for this instance. Assume that for a pair of vertices $u,v$ in $G$, there is a $(b,c)$-path gadget $P_{uv}$ present in $G$ (i.e., $P_{uv}$ is an induced subgraph of $G$) and there are more than $b$ arc-disjoint paths from $v$ to $u$. If $S$ contains an arc from $P_{uv}$, then there exists $i\in\{0,\dots,b+1\}$ such that $S$ contains every $(w_{i},w_{i+1})$ arc in $P_{uv}$.
Proof.
The argument is mostly similar to the proof of Lemma 3.1. For each $i\in[b+1]$, let $F_{i}$ be all the arcs between $w_{i-1}$ and $w_{i}$. Assume for a contradiction that we have a solution $S$ such that $S$ contains some arcs from $P_{uv}$ but there is no $i\in[b+1]$ such that $F_{i}\subseteq S$ holds, i.e., there is a path $P$ from $u$ to $v$ in $G-S$ such that all the vertices of the path belong to $P_{uv}$. As $|S|\leq b$ and $S\cap A(P_{uv})\neq\emptyset$, there is an $i\in[b+1]$ such that either $S\cap F_{i}\neq\emptyset$, $S\cap F_{i+1}=\emptyset$ or $S\cap F_{i}=\emptyset$ , $S\cap F_{i+1}\neq\emptyset$ holds. Assume that $S\cap F_{i}\neq\emptyset$, $S\cap F_{i+1}=\emptyset$. The argument in the other case is analogous. Since there is a path from $u$ to $v$ in $G-S$ and there are more than $b$ arc-disjoint paths from $v$ to $u$ in $G$, it follows that $u$ and $v$ are in the same strongly connected component in $G-S$. Moreover, since the path $P$ contains the vertices $u,w_{i-1},w_{i},w_{i+i}$ and $v$, it follows that the vertex $w_{i}$ is in the same strongly connected component in $G-S$ as $w_{i-1}$ and $w_{i+1}$. As $S\cap F_{i}\neq\emptyset$, $S\cap F_{i+1}=\emptyset$ the vertex $w_{i}$ is not balanced in $\mathrm{strong}(G-S)$, a contradiction. ∎
Brief idea of the reduction. The main idea of the following reduction is to “choose” vertices and edges of the clique using cuts. First, we enforce an imbalance using $(b,c)$-imbalance gadgets where $b$ is the budget and let it propagate using path gadgets in a way that chooses a vertex for each color. For each chosen vertex, the solution is then forced to select $(\ell-1)$ out-going arcs that are incident to it. Choosing the same edge from two sides results in a specific vertex to be cut from the strongly connected component of the remaining graph, decreasing the degree by the correct amount. Our solution creates a set of $\binom{\ell}{2}$ vertices that have out-degree two — these vertices represent edges of the multicolored clique.
See 1.1
Proof.
Consider an instance $\mathcal{I}=\big{(}G,\ell,(V_{1},\dots,V_{\ell})\big{)}$ of Multicolored Clique with $n$ vertices. Recall our assumption that each color class induces an independent set, and every vertex has at least one neighbor in every color class distinct from its own. In polynomial time, we construct an ESCAD instance $\mathcal{I}^{\prime}=(G^{\prime},k)$ in the following way (see Figure 2 for an overview).
  • We set $k=2\ell(\ell-1)$.
  • Construction of $V(G^{\prime})$ is as follows:
    1. 1.
      We add a vertex $s$.
    2. 2.
      For each color $j\in[\ell]$, we have a pair of vertices $s_{j}$ and $d_{j}$.
    3. 3.
      For each vertex $u$ in $V(G)$, we have a vertex $x_{u}$.
    4. 4.
      For each edge $uv$ in $E(G)$, we have a vertex $z_{uv}$.
  • The construction of $E(G^{\prime})$ is as follows. We introduce four sets of arcs $E_{1}$, $E_{2}$, $E_{3}$, and $E_{4}$ that together comprise the set $E(G^{\prime})$. For each color $j\in[\ell]$, let $r_{j}\coloneqq|V_{j}|\cdot(\ell-1)$, $c_{j}\coloneqq|\{uv:uv\in E(G),u\in V_{j}\}|-r_{j}$. Notice that $c_{j}\geq 0$ since every vertex in $G$ has degree at least $\ell-1$.
    1. 1.
      For each $j\in[\ell]$, we add a $(k,r_{j}-\ell+1)$-imbalance gadget $I_{s,d_{j}}$ and a $(k,c_{j})$-imbalance gadget $I_{s,s_{j}}$ to $E_{1}$.
    2. 2.
      For each $j\in[\ell]$, for each vertex $u\in V_{j}$ we add a $(k,|N_{G}(u)|-\ell+1)$-imbalance gadget $I_{s_{j},x_{u}}$ and a $(k,\ell-1)$-path gadget $P_{d_{j},x_{u}}$ to $E_{2}$.
    3. 3.
      For every edge $uv\in E(G)$, we add a pair of arcs $(x_{u},z_{uv})$ and $(x_{v},z_{uv})$ to $E_{3}$.
    4. 4.
      For every edge $uv\in E(G)$, we add two copies of the arc $(z_{uv},s)$ to $E_{4}$.
Figure 2: Overview of the reduction from $G$ to $G^{\prime}$; four sets of edges are depicted from top to bottom. $E_{1}$ contains imbalance gadgets, $E_{2}$ is a mixture of imbalance and path gadgets, $E_{3}$ has directed edges, and $E_{4}$ has all directed double edges to $s$. Marked purple edges corresponds to a solution in $G$ and its respective solution in $G^{\prime}$. The thee colored backgrounds in $G^{\prime}$ signify part of the construction tied to the three color classes. All edges of the picture of $G^{\prime}$ are oriented from top to bottom. The picture of $G^{\prime}$ wraps up as the vertex $s$ drawn on the bottom is the same as the one drawn on the top.
It is easy to see that the construction can be performed in time polynomial in $|V(G)|$. Now, we prove the correctness of our reduction. First, we argue about the imbalances of vertices in $G^{\prime}$. As each vertex of $G^{\prime}$ lies on a cycle that goes through $s$, it follows that $G^{\prime}$ is strongly connected.
Claim 3.3.
The only vertices with non-zero imbalance in $G^{\prime}$ are those in the set $\{s\}\cup\{d_{j}:j\in[\ell]\}$. Furthermore, the imbalance of the vertex $s$ is $-\ell(\ell-1)$ and the imbalance of $d_{j}$ for each $j\in[\ell]$ is $(\ell-1)$.
Proof.
There are six types of vertices in $G^{\prime}$ – (1) the vertex $s$, (2) vertices $s_{j}$ for $j\in[\ell]$, (3) vertices $d_{j}$ for $j\in[\ell]$, (4) vertices $x_{u}$ for $u\in V(G)$, (5) vertices $z_{uv}$ for $uv\in E(G)$ and (6) the intermediate vertices (in the imbalance gadgets and path gadgets). Below, we examine their imbalance one by one in the order given above.
Due to the arc set $E_{1}$, for $j\in[\ell]$, the vertex $s$ has $(k+r_{j}-\ell+1+1)+(k+c_{j}+1)=2k+r_{j}+c_{j}-\ell+3$ outgoing arcs and $(k+1)+(k+1)=2k+2$ incoming arcs. Thus, within the set $E_{1}$, each $j$ contributes $r_{j}+c_{j}-\ell+1$ more outgoing arcs than incoming arcs to $s$. So, in total the vertex $s$ has $2|E(G)|-\ell(\ell-1)$ more outgoing arcs than incoming arcs. Now, due to $E_{4}$, the vertex $s$ has $2|E(G)|$ incoming arcs. Hence, the imbalance of $s$ is $-\ell(\ell-1)$.
Due to the arc set $E_{1}$, for each $j\in[\ell]$, the vertex $s_{j}$ has $c_{j}=|\{uv:uv\in E,u\in V_{j}\}|-|V_{j}|\cdot(\ell-1)$ incoming arcs. Due to the arc set $E_{2}$, the vertex $s_{j}$ has $\sum_{u\in V_{j}}(k+|N_{G}(u)|-\ell+1+1)$ (say $q_{1}$) outgoing arcs and $\sum_{u\in V_{j}}(k+1)$ (say $q_{2}$) incoming arcs. As $c_{j}+q_{2}=q_{1}$, $s_{j}$ is balanced in $G^{\prime}$.
Within the arc set $E_{1}$, for each $j\in[\ell]$, the vertex $d_{j}$ has $r_{j}-(\ell-1)$ more incoming arcs than outgoing arcs. Due to the arc set $E_{2}$, the vertex $d_{j}$ has $|V_{j}|\cdot(\ell-1)=r_{j}$ outgoing arcs. Hence, the vertex $d_{j}$ has imbalance $(\ell-1)$.
Due to the arc set $E_{2}$, the vertex $x_{u}$ has $k+|N_{G}(u)|-\ell+1+1+\ell-1=k+|N_{G}(u)|+1$ incoming arcs and $k+1$ outgoing arcs. Due to the arc set $E_{3}$, the vertex $x_{u}$ has $|N_{G}(u)|$ outgoing arcs. Summing up, the vertex $x_{u}$ is balanced in $G^{\prime}$.
For each $uv\in E(G)$, the vertex $z_{uv}$ is incident to exactly two incoming and two outgoing arcs so it is balanced in $G^{\prime}$.
All other vertices are intermediate vertices of some imbalance or path gadget that is indeed balanced due to our construction.
This shows that there are only $\ell+1$ vertices with non-zero imbalance in $G^{\prime}$. The imbalance of the $d_{j}$’s will make us “choose” vertices and edges that represent a clique in $G$ as we will see later.
We now show correctness of our reduction. In the forward direction, assume that $(G,\ell)$ is a yes-instance and let $K$ be a multicolored clique of size $\ell$ in $G$. Let $v_{j}$ denote the vertex with color $j$ in $K$. We now construct a solution $S$ of $(G^{\prime},k)$. For each edge $v_{i}v_{j}$ we add the arcs $(x_{v_{j}},z_{v_{j}v_{i}})$ and $(x_{v_{i}},z_{v_{j}v_{i}})$ to $S$. There are 2($2\cdot\binom{\ell}{2}$ many such arcs. Now for each $j\in[\ell]$ we add all the incoming arcs of $x_{v_{j}}$ along the path gadget $P_{d_{j}x_{v_{j}}}$ to $S$. As for each $j\in[\ell]$, the number of such arcs is $(\ell-1)$ we have $|S|=2\cdot\binom{\ell}{2}+\ell(\ell-1)=2\ell(\ell-1)=k$. Now, we show that each strongly connected component in $G^{\prime}-S$ is Eulerian. For an example of $S$, refer to Figure 2 (purple arcs).
We consider the strongly connected components of $G^{\prime}-S$ and we will show that each of them is Eulerian. We first define:
$Z=\{z_{uw}:u,w\in K\}\cup\big{(}\bigcup_{j\in[\ell]}(V(P_{d_{j},x_{v_{j}}})\setminus\{d_{j},x_{v_{j}}\})\big{)}$
Claim 3.4.
One strongly connected component of $G^{\prime}-S$ consists of all the vertices except $Z$ (we call it the large component) and all other strongly connected components of $G^{\prime}-S$ are singleton – one for each vertex in $Z$.
Proof.
We have added all incoming arcs of the vertices in $\{z_{uw}:u,w\in K\}$ to $S$, so they are in singleton strongly connected components of $G^{\prime}-S$. Moreover, $x_{v_{j}}$ is a sink vertex in the path gadget $P_{d_{j}x_{v_{j}}}$. Hence, every cycle of $G^{\prime}$ in which a vertex from $(\bigcup_{j\in[\ell]}(V(P_{d_{j},x_{v_{j}}})\setminus\{d_{j},x_{v_{j}}\})$ participates, has to include an incoming arc of $x_{v_{j}}$ along the path gadget $P_{d_{j}x_{v_{j}}}$, for some $j\in[\ell]$. But these arcs have all been added to $S$. We now show that all the vertices in $G^{\prime}-S$ except those from $Z$ lie in the same strongly connected component. Consider the vertex $s$. For every $j\in[\ell]$ and $u\in V_{j}$, there is a strongly connected component in $G^{\prime}-S$ containing all the arcs from the imbalance gadgets $I_{s,s_{j}}$ (i.e., from $E_{1}$) and $I_{s_{j},x_{u}}$ (i.e., from $E_{2}$). Recall that the imbalance gadgets have arcs in both directions between consecutive vertices. Now, for any $z_{u,v}\notin Z$, we have either $u\notin K$ or $v\notin K$. Without loss of generality, assume that $u\notin K$. Then, we have a path from $x_{u}$ to $Z_{u,v}$ using an arc from $E_{3}$, so there is a path from $s$ to each $z_{u,v}\notin Z$ in $G^{\prime}-S$. In $E_{4}$, for each $z_{u,v}$ there is an arc $(z_{u,v},s)$ in $G^{\prime}-S$. Hence there is a cycle in $G^{\prime}-S$ passing through $s$ and $z_{u,v}$ for each $z_{u,v}\notin Z$. This completes the proof of the claim. ∎
Since singleton strongly connected components are always balanced, we only need to show that the large component is Eulerian i.e., it is balanced inside the strongly connected component itself.
Claim 3.5.
The large component is Eulerian
Proof.
We have the following four cases:
Consider the vertex $s$, and the large component $G^{\prime}-Z$. We have that $\deg^{+}_{G^{\prime}}(s)=\deg^{+}_{G^{\prime}-Z}(s)$ whereas the large component contains all but $\binom{\ell}{2}$ of the in-neighbors of $s$, i.e., precisely $\{z_{uw}:u,w\in K\}$. Recall that each vertex $z_{uw}$ has two arcs to $s$. So, $\deg^{-}_{G^{\prime}}(s)-\ell(\ell-1)=\deg^{-}_{G^{\prime}-Z}(s)$. As the imbalance of the vertex $s$ in $G^{\prime}$ is $-\ell(\ell-1)$ (by 3.3), the vertex $s$ is balanced in the large component.
The vertices in $\{s_{j}:j\in[\ell]\}\cup\{z_{uv}:u\notin K~{}\text{or}~{}v\notin K\}\cup\{x_{u}:u\notin K\}$ remain balanced as the large component contains all their in-neighbors and out-neighbors in $G^{\prime}$ and in $G^{\prime}$, these vertices were already balanced.
Now consider the vertex $d_{j}$ for any $j\in[\ell]$. $d_{j}$ belongs to the large component, i.e., $G^{\prime}-Z$. We have that $\deg^{-}_{G^{\prime}}(d_{j})=\deg^{-}_{G^{\prime}-Z}(d_{j})$, whereas the large component contains all but $\ell-1$ many out-neighbors of $d_{j}$, which are contained in the path gadget $P_{d_{j},v_{j}}$. So, $\deg^{+}_{G^{\prime}}(d_{j})-(\ell-1)=\deg^{+}_{G^{\prime}-Z}(d_{j})$. As the imbalance of the vertex $d_{j}$ in $G^{\prime}$ is $(\ell-1)$ (by 3.3), the vertex $d_{j}$ is balanced in the large component. By 3.3, the imbalance of the vertex $d_{j}$ in $G^{\prime}$ is $-(\ell-1)$. So $d_{j}$ is balanced in the large component.
Finally consider a vertex $x_{u}$ where $u\in K$. By 3.3, this vertex is balanced in $G^{\prime}$. But the large component does not contain all the neighbors of $x_{u}$. It excludes $\ell-1$ out-neighbors which are precisely $\{z_{uv}:v\in K\}$ and $\ell-1$ in-neighbors which come from precisely one path gadget $P_{d_{j},x_{u}}$ where $u\in V_{j}$. So, this vertex is also balanced in the large component.
This completes the proof of the claim. ∎
This completes the argument in the forward direction.
In the converse direction, assume that $(G^{\prime},k)$ is a yes-instance and let $S$ be a solution. Let us first establish some structure on $S$, from which it will be possible to recover a multicolored clique for $G$.
Let $\mathcal{C}$ denote the strongly connected component of $G^{\prime}-S$ that contains $s$. Due to Lemma 3.1, we may assume that $S$ does not contain any arcs of any of the imbalance gadgets. This implies that $\mathcal{C}$ contains $s_{j}$ and $d_{j}$ for every $j\in[\ell]$ as well as $x_{u}$ for every $u\in V(G)$. Moreover, due to Lemma 3.2, we know that if $S$ contains arcs of a path gadget $P_{d_{j},x_{u}}$, then they form a cut in it. As all inclusion-wise minimal cuts of the path gadgets are of the same cardinality and adding any minimal cut of a path gadget to $S$ makes all arcs of the path gadget inactive in $G^{\prime}-S$, assume that if a cut of a path gadget $P_{d_{j},x_{u}}$ is in $S$, then the cut consists of the incoming-arcs of $x_{u}$ in the gadget.
Recall from 3.3, that the only imbalanced vertices in $G^{\prime}$ are $\{s\}\cup\{d_{j}:j\in[\ell]\}$. Let us make some observations based on the fact that these vertices are eventually balanced in $\mathrm{strong}(G^{\prime}-S)$.
For each $j\in[\ell]$, since none of the incoming arcs of $d_{j}$ are in $S$ (they lie in an imbalance gadget), in order to make $d_{j}$ balanced it must be the case that $S$ contains a cut of exactly one of the path gadgets starting at $d_{j}$, call it $P_{d_{j},x_{v_{j}}}$. Recall that $x_{v_{j}}$ was originally balanced in $G^{\prime}$. Further, recall that we have argued that $x_{v_{j}}$ is in $\mathcal{C}$ along with $s_{j}$ and $d_{j}$. Since the imbalance gadget starting at $s_{j}$ and ending at $x_{v_{j}}$ cannot intersect $S$ and we have deleted all of the $\ell-1$ incoming arcs to $x$ from the path gadget $P_{d_{j},x_{v_{j}}}$, the imbalance of $-\ell+1$ thus created at $x_{v_{j}}$ needs to be resolved by making exactly $(\ell-1)$ of its outgoing arcs in $E_{3}$ inactive in $G^{\prime}-S$. Since we have already spent a budget of $\ell(\ell-1)$ from the path gadgets, the budget that remains to be used for resolving these imbalances at $\{x_{v_{j}}:j\in[\ell]\}$ is $\ell(\ell-1)$.
On the other hand, recall that $s$ is imbalanced in $G^{\prime}$ and to make $s$ balanced, we need to make $\ell(\ell-1)$ incoming arcs of $s$ (from $E_{4}$) inactive in $G^{\prime}-S$. This is because all outgoing arcs of $s$ lie in imbalance gadgets and cannot be in $S$.
And finally, recall that for each $uv\in E(G)$, the vertex $z_{uv}$ is balanced in $G^{\prime}$ (by 3.3). Since the strongly connected component $\mathcal{C}$ in $G^{\prime}-S$ contains the vertices $s,x_{u},x_{v}$ (i.e., all neighbors of $z_{u,v}$), for the vertex $z_{uv}$ to remain balanced in $\mathrm{strong}(G^{\prime}-S)$, we have the following exhaustive cases regarding the arcs between $s,x_{u},x_{v},z_{u,v}$: (1) none of the four arcs incident to $z_{uv}$ is in $S$; (2) one incoming and one outgoing arc are in $S$; (3) both incoming arcs or both outgoing arcs are in $S$. In Case (2), two arcs are added to $S$, which makes two arcs inactive while in Case (3) two arcs are added to $S$ which makes four arcs inactive. As previously noted, we still need $\ell(\ell-1)$ arcs in $E_{3}$ and $\ell(\ell-1)$ arcs in $E_{4}$ to become inactive in $G^{\prime}-S$. The required number of inactive arcs in $E_{3}\cup E_{4}$ is twice the remaining budget, so for every $z_{uv},x_{u},x_{v}$, the arcs between $s,x_{u},x_{v},z_{u,v}$ must be in Case (1) or Case (3). Moreover, whenever Case (3) occurs, we may assume without loss of generality that the arcs in $S$ are the two arcs $(x_{u},z_{uv})$ and $(x_{v},z_{uv})$. Thus, there are exactly $\binom{\ell}{2}$ vertices $z_{uv}$ such that the arcs between $s,x_{u},x_{v},z_{u,v}$ are in Case (3).
We now extract the solution clique $K$ for $(G,\ell)$ by taking, for each $j\in[\ell]$, the vertex $v_{j}\in V(G)$ such that a cut of $P_{d_{j},x_{v_{j}}}$ is contained in $S$. We have shown that there are exactly $\binom{\ell}{2}$ vertices $z_{uv}$ such that the arcs between $s,x_{u},x_{v},z_{u,v}$ are in Case (3) and for each $j\in[\ell]$ and the vertex $x_{v_{j}}$, exactly $\ell-1$ of its outgoing arcs are made inactive by $S$. This can only happen if for every $j,j^{\prime}\in[\ell]$, there is a vertex $z_{v_{j}v_{j^{\prime}}}$, implying that $v_{j}v_{j^{\prime}}$ is an edge in $G$. ∎
3.2 W[1]-hardness of ESCAD Parameterized by Vertex Cover Number
In this section, we show that ESCAD is W$[1]$-hard when parameterized by the vertex cover number. Jansen, Kratsch, Marx, and Schlotter [12] showed that Unary Bin Packing is W$[1]$-hard when parameterized by the number of bins $h$.
Unary Bin Packing Input: A set of positive integer item sizes $x_{1},\dots,x_{n}$ encoded in unary, a pair of integers $h$ and $b$. Question: Is there a partition of $[n]$ into $h$ sets $J_{1},\dots,J_{h}$ such that $\sum_{\ell\in J_{j}}x_{\ell}\leq b$ for every $j\in[h]$?
In order to carefully handle vertex balances in our reduction, it is helpful to work with a variant of the above problem, called Exact Unary Bin Packing, where the inequality $\sum_{\ell\in J_{j}}x_{\ell}\leq b$ is replaced with the equality $\sum_{\ell\in J_{j}}x_{\ell}=b$. That is, in this variant, all bins get filled up to their capacity.
See 1.2
Proof.
Let $\mathcal{I}^{\prime}=\big{(}(x_{1},\dots,x_{m}),h,b\big{)}$ be an instance of Unary Bin Packing. If $b\geq\sum_{i=1}^{m}x_{i}$, then $\mathcal{I}^{\prime}$ is trivially a yes-instance and we can return a trivial yes-instance of ESCAD with vertex cover number at most $h$. In the same way, if $b\cdot h<\sum_{i\in[m]}x_{i}$, then $\mathcal{I}^{\prime}$ is trivially a no-instance and we return a trivial no-instance of ESCAD with vertex cover number at most $h$. Now, suppose neither of the above cases occur.
Note that the length of the unary encoding of $b$ is upper bounded by the total length of the unary encoding of all items $x_{1},\dots,x_{m}$. Similarly, if $h\geq m$ then the instance boils down to checking whether $x_{i}\leq b$ for every $i\in[m]$ (and producing a trivial ESCAD instance accordingly) so we can assume that $h<m$, hence, the length of the unary encoding of $h$ is upper bounded by the total length of the unary encoding of all items. We now construct an instance $\mathcal{I}$ of Exact Unary Bin Packing from $\mathcal{I}^{\prime}$ by adding $h\cdot b-\sum_{i\in[m]}x_{i}$ one-sized items (this is non-negative because of the preprocessing steps). If $\mathcal{I}^{\prime}$ is a yes-instance, then one can fill-in the remaining capacity in every bin with the unit-size items, to get a solution for $\mathcal{I}$. Conversely, if $\mathcal{I}$ is a yes-instance, then removing the newly added unit-size items yields a solution for $\mathcal{I}^{\prime}$. Let $n$ denote the number of items in $\mathcal{I}$. Note that since $\sum_{i\in[n]}x_{i}=b\cdot h$, this implies that $|\mathcal{I}|=\mathcal{O}(|\mathcal{I}^{\prime}|^{2})$, the instance of Exact Unary Bin Packing remains polynomially bounded.
We next reduce the Exact Unary Bin Packing instance $\mathcal{I}$ to an instance $\mathcal{I}^{*}=(G,k)$ of ESCAD in polynomial time. Let us fix the budget $k=b\cdot h(h-1)$. We now build a graph $G$ that models the bins by $k$ copies of interconnected gadgets (that form the vertex cover) and models each item as a vertex of the independent set. In our reduction, we use the following terms. For a pair of vertices $p,q$, a $c$-arc $(p,q)$ denotes $c$ parallel copies of the arc $(p,q)$ and a thick arc $(p,q)$ denotes a $3k$-arc $(p,q)$. The construction of $G$ is as follows.
  • The vertex set of $G$ is the set $\{u_{j}:j\in[h]\}\cup\{v_{j}:j\in[h]\}\cup\{w_{i}:i\in[n]\}$.
  • For each $j\in[h]$, we add a $b$-arc $(u_{j},v_{j})$, a thick arc $(u_{j},v_{j})$ and a thick arc $(v_{j},u_{j})$. We call the subgraph induced by $u_{j},v_{j}$ and these arcs, the $b$-imbalance gadget $B_{j}$.
  • Next, we add thick arcs $(u_{j},u_{j^{\prime}})$ for every $j<j^{\prime}$ where $j,j^{\prime}\in[h]$.
  • Finally, for each $i\in[n]$ and $j\in[h]$, we add $x_{i}$-arcs $(w_{i},u_{j})$ and $(v_{j},w_{i})$.
This concludes the construction, see Figure 3. Before we argue the correctness, let us make some observations.
Note that the vertices participating in the imbalance gadgets form a vertex cover of the resulting graph and their number is upper bounded by $2h$. Hence, if we prove the correctness of the reduction, we have the required parameterized reduction from Unary Bin Packing parameterized by the number of bins to ESCAD parameterized by the vertex cover number of the graph.
We say that a set of arcs in $G$ cuts a $(p,q)$ arc if it contains all parallel copies of $(p,q)$. Note that no set of at most $k$ arcs cuts a thick $(p,q)$ arc. In particular, no solution to the ESCAD instance $(G,k)$ cuts any thick arc $(p,q)$ that appears in the graph.
Figure 3: A part of the resulting ESCAD instance after reduction from Exact Unary Bin Packing with six bins; connections between the independent vertices and imbalance gadgets are shown only for one vertex $w_{i}$. Thick arcs are shown with empty arrowhead, bold arcs incident to $w_{i}$ are $x_{i}$-arcs. Crossed off arcs are in a solution and dashed boxes show strongly connected components of the solution. This example represents $x_{i}\in J_{4}$.
Exact Unary Bin Packing is a yes-instance $\Rightarrow$ ESCAD is a yes-instance: Assume that we have a partition $J_{1},\dots,J_{h}$ that is a solution to $\mathcal{I}$. We now define a solution $S$ for $\mathcal{I}^{*}$. For every $x_{i}\in J_{j}$ we cut (i.e., add to $S$) all parallel copies of the arc $(w_{i},u_{j^{\prime}})$ for every $j^{\prime}<j$ and we cut all parallel copies of the arc $(v_{j^{\prime\prime}},w_{i})$ for every $j^{\prime\prime}>j$. This results in cutting a total of $x_{i}\cdot(h-1)$ arcs incident to each $w_{i}$ and as $\sum_{i=1}^{n}x_{i}=b\cdot h$ we cut exactly $b\cdot h(h-1)=k$ arcs in total.
Claim 3.6.
$\mathrm{strong}(G-S)$ is balanced.
Proof.
Due to the thick $(u_{j},v_{j})$ and $(v_{j},u_{j})$ arcs and our earlier observation that no set of at most $k$ arcs can cut a thick arc, we have that for every $j\in[h]$, there is a single strongly connected component of $G-S$ containing both vertices of $B_{j}$. We next observe that in $\mathrm{strong}(G-S)$ no pair of distinct $b$-imbalance gadgets are contained in the same strongly connected component. This is because any path in $G$ from $B_{j}$ to $B_{i}$ for $i<j$ must use arcs $(v_{j^{\prime}},w_{p})$ and $(w_{p},u_{i^{\prime}})$ for some $p\in[n]$, and $i^{\prime},j^{\prime}\in[h]$ such that $i^{\prime}<j^{\prime}$. However, one of the these two arcs is part of $S$ by definition.
Further, notice that the strongly connected component containing $B_{j}$ also contains the vertex $w_{i}$ if $x_{i}\in J_{j}$. This is because we do not delete the arcs $(v_{j},w_{i})$ and $(w_{i},u_{j})$. Since we have already argued that the imbalance gadgets are all in distinct strongly connected components in $G-S$, we infer that the strongly connected component containing $B_{j}$ also contains the vertex $w_{i}$ if and only if $x_{i}\in J_{j}$. Hence, we conclude that incident to $w_{i}$, the only active arcs are those of the form $(w_{i},u_{j})$ and $(v_{j},w_{i})$ for $j$ such that $x_{i}\in J_{j}$, making $w_{i}$ balanced in $\mathrm{strong}(G-S)$. For $u_{j}$ and $v_{j}$ for $j\in[h]$, the thick $(u_{j},v_{j})$ and thick $(v_{j},u_{j})$ arcs balance each other. The only active arcs that remain and are incident on $u_{j}$ and $v_{j}$ are the $b$-arcs $(u_{j},v_{j})$ for each $j\in[h]$. We argue that these are balanced by the arcs incoming from all the vertices $w_{i}$ to $u_{j}$ and the arcs outgoing from $v_{j}$ to $w_{i}$, where $x_{i}\in J_{j}$. Indeed, $\sum_{\ell\in J_{j}}{x_{\ell}}=b$ for all $j\in[h]$ so $u_{j}$ has $b$ incoming arcs and $v_{j}$ has $b$ outgoing arcs from the vertices $\{w_{1},\dots,w_{n}\}$, making $u_{j}$ and $v_{j}$ balanced in $\mathrm{strong}(G-S)$ for all $j\in[h]$. ∎
ESCAD is a yes-instance $\Rightarrow$ Exact Unary Bin Packing is a yes-instance: We aim to show that in any solution for the ESCAD instance, the arcs that are cut incident to $w_{i}$ for any $i\in[n]$ have the same structure as described in the other direction, i.e., for all $w_{i}$ there exists $j$ such that the solution cuts $(w_{i},u_{j^{\prime}})$ for all $j^{\prime}<j$ and it cuts $(v_{j^{\prime\prime}},w_{i})$ for all $j^{\prime\prime}>j$. This is equivalently phrased in the following claim.
Claim 3.7.
There are no two indices $a,b\in[h]$ with $a<b$ such that both $(w_{i},u_{a})$ and $(v_{b},w_{i})$ are uncut.
Proof.
Towards a contradiction, consider a solution $S$ without this property. That is, for some $1\leq a<b\leq h$, both $(w_{i},u_{a})$ and $(v_{b},w_{i})$ are uncut by $S$. The graph $G$ contains thick arcs $(u_{a},u_{b})$ and $(u_{b},v_{b})$ that cannot be cut by $S$. Hence, there is a cycle $(w_{i},u_{a},u_{b},v_{b},w_{i})$ in $\mathrm{strong}(G-S)$, implying that the vertices of two imbalance gadgets $B_{a}$ and $B_{b}$ are in the same strongly connected component of $G-S$. Choose $a,b$ such that $a$ is minimized. We argue that $u_{a}$ cannot be balanced in $\mathrm{strong}(G-S)$. We first ignore the two thick arcs $(u_{a},v_{a})$ and $(v_{a},u_{a})$ as they balance each other. We picked $a$ to be minimum, so $u_{a}$ has no active incoming arcs that belong to a thick $(u_{a^{\prime}},u_{a})$ arc for some $a^{\prime}<a$ since otherwise, $u_{a},v_{a},u_{a^{\prime}},v_{a^{\prime}}$ would be in the same strongly connected component of $G-S$. Hence, the only remaining active incoming arcs on $u_{a}$ are the incoming arcs from $\{w_{1},\dots,w_{n}\}$, of which there are exactly $\sum_{i\in[n]}x_{i}$ arcs. Recall that we have $\sum_{i\in[n]}x_{i}=b\cdot h$ and by the definition of $k$, we have $h\cdot b<k$. This implies that in $G-S$, $u_{a}$ has at least $2k$ active outgoing arcs (at most $k$ out of the $3k$ arcs contained in the thick $(u_{a},u_{b})$ arc can be in $S$) and at most $k$ incoming active arcs, a contradiction to $S$ being a solution. Hence, we conclude that for all $w_{i}$ there exists $j$ such that the solution cuts $(w_{i},u_{j^{\prime}})$ for all $j^{\prime}<j$ and it cuts $(v_{j^{\prime\prime}},w_{i})$ for all $j^{\prime\prime}>j$. ∎
We next argue that if $S$ is a solution, then for all $w_{i}$, there exists $j$ such that the solution is disjoint from any $(w_{i},u_{j})$ arc and any $(v_{j},w_{i})$ arc. Since the budget is $k=b\cdot h(h-1)$ we have that: If we cut more than $x_{i}(h-1)$ arcs incident to $w_{i}$ for some $i\in[n]$, then there exists $i^{\prime}\in[n]\setminus\{i\}$ such that we cut fewer than $x_{i^{\prime}}(h-1)$ arcs incident to $w_{i^{\prime}}$. But this would violate 3.7. Hence, for any solution $S$, we can retrieve the assignment of items to bins in the Exact Unary Bin Packing instance $\cal I$, by identifying for every $i\in[n]$, the unique value of $j\in[h]$ such that $S$ is disjoint from any $(w_{i},u_{j})$ arc and any $(v_{j},w_{i})$ arc and then assigning item $x_{i}$ to bin $J_{j}$. ∎
Besides establishing that Unary Bin Packing does not have an FPT algorithm unless W$[1]$ = FPT, Jansen et al. [12] showed that under the stronger assumption [^3] It is known that if ETH is true, then W[1]delimited-[]1absent[1$\neq$[ 1 ] ≠ FPT [6]. of the Exponential Time Hypothesis (ETH) the well-known $n^{{\mathcal{O}}(h)}$-time algorithm is asymptotically almost optimal The formal statement follows.
Proposition 3.8 ([12]).
There is no algorithm solving the Unary Bin Packing problem in $f(h)\cdot n^{o(h/\log h)}$ time for some function $f$, where $h$ is the number of bins in the input and $n$ is the input length, unless ETH fails.
Since our reduction from Unary Bin Packing to ESCAD transforms the parameter linearly and the instance size polynomially, we also have a similar ETH based lower bound parameterized by the vertex cover number for ESCAD.
See 1.3
Proof.
Follows from the reduction in the proof of Theorem 1.2 along with Proposition 3.8. ∎
3.3 NP-hardness of ESCAD on Graphs of Constant Maximum Degree
We show that ESCAD is para-NP-hard when parameterized by the maximum degree.
See 1.4
Proof.
We give a polynomial-time reduction from Vertex Cover on cubic (3-regular) graphs, which is known to be NP-hard [15], to ESCAD. This reduction is a modification of the proof in [15] which shows that Directed Feedback Arc Set is NP-hard. The input to Vertex Cover consists of a graph $G$ and an integer $k$; the task is to decide whether $G$ has a vertex cover of size at most $k$. Let $(G,k)$ be an instance of Vertex Cover with $n$ vertices where $G$ is a cubic graph. We construct an ESCAD instance $\mathcal{I}^{\prime}=(G^{\prime},k)$ in the following way. The vertex set $V(G^{\prime})=V(G)\times\{0,1\}$ and the arc set $A(G^{\prime})$ is defined by the union of the sets $\{((u,0),(u,1)):u\in V(G)\}$ and $\{((u,1),(v,0))^{2}:uv\in E(G)\}$. We call the arcs of the form $((u,0),(u,1))$ internal arcs and arcs of the form $((u,1),(v,0))$ cross arcs. Towards the correctness of the reduction, we prove the following claim.
Claim 3.9.
$(G,k)$ is a yes-instance of Vertex Cover if and only if $(G^{\prime},k)$ is a yes-instance of ESCAD.
Proof.
In the forward direction, let $(G,k)$ be a yes-instance and let $X$ be a solution. Consider the arc set $F=\{((u,0),(u,1)):u\in X\}\subseteq A(G^{\prime})$. We show that $F$ is a feedback arc set of $G^{\prime}$. Consider any cycle in $G^{\prime}$. Due to our construction, the cycle must two internal arcs $((u,0),(u,1))$ and $((v,0),(v,1))$ where $uv\in E(G)$. Now either $u\in X$ or $v\in X$. That implies that either $((u,0),(u,1))$ or $((v,0),(v,1))$ belongs to $F$. Hence $G^{\prime}-F$ has no cycles. As $G^{\prime}-F$ is acyclic, we have that $F$ is a solution to the ESCAD instance $(G^{\prime},k)$.
In the converse direction, let $(G^{\prime},k)$ be a yes-instance, let $F$ be a solution for this instance with minimum number of cross arcs. We first argue that $G^{\prime}-F$ is acyclic. Suppose not. Because of the structure of the constructed digraph, every cycle alternates between internal and cross arcs. So, every strongly connected component in $G^{\prime}-F$ must contain an internal arc, and as it must also be Eulerian, the strongly connected component must be a simple cycle $C$ (as for each $u\in V(G)$, the out-degree of the vertex $(u,0)$ in $G^{\prime}-F$ is at most one). Each arc of $C$ is present only once – to achieve that, the solution $F$ must contain at least one copy of each of the cross arcs that are in $C$. Now, we can remove all the copies of cross arcs in $C$ from the solution and instead, add all internal arcs of $C$ to the solution. This gives us a new solution with fewer cross arcs, a contradiction to our choice of $F$. Hence, we may assume that $G^{\prime}-F$ is acyclic. We now argue that $X=\{u:((u,0),(u,1))\in F\}\cup\{u:((u,1),(v,0))\in F\}$ is a vertex cover of $G$ of size at most $k$. Clearly $|X|\leq k$. Consider an arbitrary edge $uv\in E(G)$. Corresponding to the edge $uv$ there is a 4-cycle $((u,0),(u,1))$, $((u,1),(v,0)),$ $((v,0),(v,1))$, $((v,1),(u,0))$ in $G^{\prime}$, and so, $F$ must contain one of these four arcs. Now, by our definition of $X$, $X\cap\{u,v\}\neq\emptyset$, hence $X$ is a solution for the Vertex Cover instance $(G,k)$. ∎
This shows that ESCAD is NP-hard. Moreover, Since $G$ is a cubic graph, every vertex in $D^{\prime}$ has (in, out) degree equal to $(1,6)$ or $(6,1)$. This completes the proof of Theorem 1.4. ∎
3.4 Algorithms for ESCAD on Graphs of Bounded Treewidth
Due to Theorem 1.2, the existence of an FPT algorithm for ESCAD parameterized by various width measures such as treewidth is unlikely. In fact, due to Theorem 1.3, assuming ETH, even obtaining an algorithm with running time $f(k)n^{o(k/\log k)}$ is not possible, where $k$ is the vertex cover number. On the other hand, this raises a natural algorithmic question – could one obtain an algorithm whose running time matches this lower bound? In this section, we give such an algorithm that is simultaneously, an XP algorithm parameterized by treewidth, an FPT algorithm parameterized by the treewidth and solution size, and also an FPT algorithm parameterized by the treewidth and maximum degree of the input digraph. Moreover, the running time of the algorithm nearly matches the lower bound we have.
Let us note that in the specific case of parameterizing by treewidth and maximum degree, if all we wanted was an FPT algorithm, then we could use Courcelle’s theorem at the cost of a suboptimal running time. However, our algorithm in one shot gives us three consequences and as stated earlier, achieves nearly optimal dependence on the treewidth assuming ETH.
Overview of our algorithm. We present a dynamic programming algorithm over tree decompositions. When one attempts to take the standard approach, the main challenge that arises is that by disconnecting strongly connected components, removing an arc can affect vertices far away and hence possibly vertices that have already been forgotten at the current stage of the algorithm. Our solution is to guess the partition of each bag into strongly connected components in the final solution and then keep track of the imbalances of the vertices of the bag under this assumption of components. This allows us to safely forget a vertex as long as its “active” imbalance is zero (any remaining imbalance will be addressed by not strongly connecting the contributing vertices in the future). The remaining difficulty lies in keeping track of how these assumed connections interact with the bag: whether they use vertices already forgotten or those yet to be introduced.
See 1.5
Since the maximum degree is upper bounded by the instance length (recall footnote in Section 1), this gives an XP algorithm parameterized by treewidth alone. However, when in addition to treewidth we parameterize either by the size of the solution or by the maximum degree this gives an FPT algorithm.
Corollary 3.10.
ESCAD is FPT parameterized by ${\mathtt{tw}}+k$, FPT parameterized by ${\mathtt{tw}}+\Delta$, and XP parameterized by ${\mathtt{tw}}$ alone.
Recall that in digraphs, multiarcs are permitted. So, we use a variant of the nice tree decomposition notion. This is defined for a digraph $G$ by taking a nice tree decomposition with introduce edge nodes (see Section 2) of the deoriented, simple version of $G$ then expanding each introduce edge node to introduce all parallel copies of arcs one by one. Note that although the new introduce arc nodes introduce arcs, the orientation does not affect the decomposition. Let us denote such a tree decomposition of $G$ as $(\mathcal{T},\{X_{t}\}_{t\in V(T)})$. Korhonen and Lokshtanov [16] gave a $2^{tw^{2}}\cdot n^{{\mathcal{O}}(1)}$-time algorithm that computes an optimal tree decomposition. Moreover, any tree decomposition can be converted to a nice tree decomposition of the same width with introduce edge nodes in polynomial time [4], and the introduce edge nodes can clearly be expanded to introduce arc nodes in polynomial time. Since the running time of our algorithm dominates the time taken for this step, we may assume that we are given such a tree decomposition. Let $G_{t}$ be the subgraph of the input graph that contains the vertices and arcs introduced in the subtree rooted at $t$. We refer to $G_{t}$ as the past and to all other arcs and vertices as the future.
To tackle ESCAD we need to know whether an arc between vertices in a bag is active in the graph minus a hypothetical solution or not. Towards this, we express the reachability of the graph that lies outside (both past and future) of the current bag as follows.
Definition 3.11.
For a set $X$, let $(R,\ell)$ be a reachability arrangement on $X$ where $R$ is a simple digraph with $V(R)=X$, and $\ell$ is a labeling $\ell\colon E(R)\to\{\mathrm{direct},\mathrm{past},\mathrm{future}\}$.
Let us use $\ell(u,v)$ to denote $\ell((u,v))$. As reachability arrangement implies which vertices of the bag lie in the same strongly connected components we can determine whether an arc is active by checking that its endpoints lie in the same strongly connected component. We aim to track the balance of the vertices in the bag with respect to all past active arcs.
Definition 3.12.
Given $G$ and $R$ the active imbalance $b_{G}^{R}(v)$ of a vertex $v$ in $G$ with respect to $R$ is the imbalance of $v$ in the graph $H$, i.e. $\deg^{+}_{H}(v)-\deg^{-}_{H}(v)$, where $H$ is the graph induced on $G$ by the vertices of the strongly connected component of $R$ containing $v$.
Although the active imbalance is bounded by $\Delta$, it can be large even when the solution is bounded so we want to instead track how much the active imbalance varies between two graphs.
Definition 3.13.
Given $G_{1}$, $G_{2}$, and $R$ the offset imbalance of a vertex $v$ between $G_{1}$ and $G_{2}$ with respect to $R$, ${\mathrm{off}}_{G_{1},G_{2}}^{R}(v)=b_{G_{1}}^{R}(v)-b_{G_{2}}^{R}(v)$.
We will consider the offset imbalance between $G_{t}$ and $G_{t}-S$ where $S$ is part of a solution. The following lemma allows us to bound this quantity by the size of the solution.
Lemma 3.14.
For each set of arcs $S\subseteq E(G)$, node $t\in V(\mathcal{T})$, simple digraph $R$ on $X_{t}$ and vertex $v\in X_{t}$, the offset imbalance of $v$ between $G_{t}-S$ and $G_{t}$ with respect to $R$ is between $-|S|$ and $|S|$.
Proof.
We have ${\mathrm{off}}_{G_{t}-S,G_{t}}^{R}(v)=b_{G_{t}-S}^{R}(v)-b_{G_{t}}^{R}(v)$. Both its terms $b_{G_{t}}^{R}(v)$ and $b_{G_{t}-S}^{R}(v)$ are calculated with respect to $R$ so the considered strongly connected components containing $v$ are the same. Let us denote vertices of the considered strongly connected component by $H$. Observe, that $b_{G_{t}}^{R}(v)$ is the difference of in-degree and out-degree of $v$ in $G_{t}[H]$. Similarly, $b_{G_{t}-S}^{R}(v)$ is the difference of in-degree and out-degree of $v$ in $(G_{t}-S)[H]$. To get from $G_{t}[H]$ to $(G_{t}-S)[H]$ we remove arcs of $S$ from $G_{t}$ one by one and note that each removal changes the degrees of its endpoints by at most one. After considering all arcs of $S$ the value of $b_{G_{t}}^{R}(v)$ could have changed by at most $|S|$ so we have $-|S|\leq{\mathrm{off}}_{G_{t},G_{t}-S}^{R}(v)\leq|S|$. ∎
For a solution $S$ we use a suitable reachability arrangement $(R,\ell)$, balance labeling $B$, and part of the solution in the bag $W$ to express a partial solution, that is: $S\cap G_{t}$ along with how vertices of the bag are partitioned into strongly connected components in $G-S$. These give a description of partial solutions that is small enough to guess but detailed enough to admit a dynamic programming approach.
Definition 3.15.
Given a node of the tree decomposition $t$, a reachability arrangement $(R,\ell)$ on $X_{t}$, a labeling $b\colon V(R)\to[-\alpha,\alpha]$, and a subset of arcs $W\subseteq E(G_{t}[X_{t}])$ we call a set of arcs $S\subseteq E(G_{t})$ compatible with $R,\ell,b,W$ if all of the following parts hold.
  1. 1.
    $S$ agrees with $W$ on $G_{t}[X_{t}]$, that is $S\cap E(G_{t}[X_{t}])=W$.
  2. 2.
    For each arc $e\in\ell^{-1}(\mathrm{direct})$, $e$ is an arc in $G_{t}[X_{t}]-S$.
  3. 3.
    For each arc $(u,w)\in\ell^{-1}(\mathrm{past})$ there is a path from $u$ to $w$ in $G_{t}-S$ that contains no vertices from $X_{t}\setminus\{u,w\}$ (also called path through the past).
  4. 4.
    For each arc $(u,w)\in\ell^{-1}(\mathrm{future})$ there is no path through the past from $u$ to $w$ (see part 3) and there is a path from $u$ to $w$ in $G-S$ that contains no vertices from $X_{t}\setminus\{u,w\}$ (also called path through the future).
  5. 5.
    For each vertex $u\in X_{t}$, the offset imbalance of $u$ between $G_{t}-S$ and $G_{t}$ with respect to $R$ is $b(u)$, i.e., ${\mathrm{off}}_{G_{t},G_{t}-S}^{R}(u)=b(u)$.
  6. 6.
    For each vertex $u\in V(G_{t})\setminus X_{t}$, the active imbalance of $u$ in $G_{t}-S$ with respect to $(G_{t}-S)\cup R$ is zero, i.e., $b_{G_{t}-S}^{(G_{t}-S)\cup R}(u)=0$.
Observation 3.16.
Suppose that $S$ is a solution. For all nodes $t$ there exists $R,\ell,b,W$ such that $S_{1}=S\cap E(G_{t})$ is compatible with $R,\ell,b,W$.
Proof.
We just note how to create the sets as it is straight-forward to check $S_{1}$ is compatible with them. Set $R$ contains an arc $(u,v)$ if and only if there exists a path from $u$ to $v$ in $G-S$. We set label $\ell$ of $(u,v)$ to be $\mathrm{direct}$ if $(u,v)\in E(G_{t})$, to $\mathrm{past}$ if there exists a path from $u$ to $v$ through the past, and otherwise we set it to $\mathrm{future}$. We compute $b(u)$ for $u\in X_{t}$ by first computing the strongly connected components of $G-S$ and then computing active degrees $b_{G_{t}}^{G-S}(u)$ and $b_{G_{t}-S}^{G-S}(v)$, setting $b(u)$ to be their difference. Since $S_{1}\subseteq S$, this $b$ has a range of $[-\alpha,\alpha]$ by Lemma 3.14 and the observation that active imbalance (and hence offset imbalance) is bounded by $\Delta$. We set $W=G_{t}[X_{t}]\cap S$. ∎
Lemma 3.17.
Suppose that $S$ is a solution and both $S_{1}=S\cap E(G_{t})$ and $S_{2}\subseteq E(G_{t})$ are compatible with $R,\ell,b,W$. Then $S^{\prime}=(S\setminus S_{1})\cup S_{2}$ is also a solution.
Proof.
It suffices to show that the active imbalance of all vertices $v\in V(G)$ with respect to $G-S^{\prime}$ is zero which we prove using parts from Definition 3.15. By part 1 we have $E(G_{t}[X_{t}])\cap S_{1}=E(G_{t}[X_{t}])\cap S_{2}$ so $S_{2}$ differs from $S_{1}$ only in $G_{t}-X_{t}$. Both $S_{1}$ and $S_{2}$ are compatible with $(R,\ell)$ so for by part 3 for any $u,v\in X_{t}$ we have that $S_{2}$ cuts a path from $u$ to $v$ through the past if and only if $S_{1}$ cuts the path. Hence, the connectivity between vertices of $X_{t}$ in $G-S^{\prime}$ is the same as in $G-S$, in fact it is exactly $R$ (by parts 2, 3, and 4). Moreover, the set of active arcs incident to vertices $G\setminus G_{t}$ in $G-S$ is the same as in $G-S^{\prime}$, implying that the active imbalance of all vertices in $G\setminus G_{t}$ is still zero. Active imbalance of vertices in $X_{t}$ comprises of active arcs to $G\setminus G_{t}$ and active arcs to $G_{t}$. We just saw that the active arcs to $G\setminus G_{t}$ do not change and by part 5 we know that the active imbalance of vertices of the bag in $G_{t}-S^{\prime}$ with respect to $R$ is the same as that in $G_{t}-S$ with respect to $R$ so the active imbalance of vertices of $X_{t}$ in $G-S^{\prime}$ remains zero. Finally, part 6 ensures that the vertices of $G_{t}-X_{t}$ have active imbalance zero in $G_{t}-S^{\prime}$ with respect to $(G_{t}-S^{\prime})\cup R$. This imbalance remains zero also in $G-S^{\prime}$ with respect to $G-S^{\prime}$ because arcs of $G$ that are incident to $G\setminus G_{t}$ are not incident to the vertices in $G_{t}-X_{t}$. ∎
The above lemma implies that for fixed $t,R,\ell,b,W$ all solutions $S$ have the same cardinality of $S\cap G_{t}$. For fixed $t,R,\ell,b,W$ to compute existence of some solution $S$ such that $S_{1}=S\cap G_{t}$ is compatible with $R,\ell,b,W$, it suffices to compute the minimum cardinality of a subset $S_{2}\subseteq E(G_{t})$ compatible with $R,\ell,b,W$ because one can always produce the solution $S^{\prime}=(S-S_{1})\cup S_{2}$.
Proof of Theorem 1.5.
We will denote by $A[t,R,\ell,b,W]$ the minimum size of an arc subset of $G_{t}$ that is compatible with $R$, $\ell$, $b$, and $W$. In our decomposition $(\mathcal{T},\{X_{t}\}_{t\in V(T)})$ the root node $r$ has $X_{r}=\emptyset$ and $G_{r}=G$ so $A[r,\emptyset,\emptyset,\emptyset,\emptyset]$ is equal to the minimum size of a solution. In order to compute $A[r,\emptyset,\emptyset,\emptyset,\emptyset]$ we employ the standard bottom up dynamic programming over treewidth decomposition approach.
For leaf nodes $X_{t}=\emptyset$, hence, the graphs and labelings are also empty and the empty arc set is vacuously compatible with them $A[t,\emptyset,\emptyset,\emptyset,\emptyset]=0$.
For every non-leaf node $t$ and graph $R$ on $X_{t}$ we first calculate the strongly connected components of $R$. Then we can calculate the active imbalance $b_{G_{t}}^{R}(v)$ of each vertex $v\in X_{t}$ in $G_{t}$ with respect to $R$. Then for each $\ell$, $b$, and $W$ we calculate $A[t,R,\ell,b,W]$ based on the type of the node $t$.
Introduce vertex node:
When $t$ is an introduce vertex node and its child is $t^{\prime}$ with $X_{t}=X_{t^{\prime}}\cup\{v\}$ we know that $v$ will be isolated in $G_{t}$ so we can discount any reachability arrangements where there are direct or past arcs incident to $v$. Additionally, the active imbalance on $v$ must be zero. Any new future connections should be reflected in the old reachability arrangement, that is, if the new arrangement contains a future arc from $u$ to $v$ and from $v$ to $w$ there should be a future arc between $u$ and $w$ in the old arrangement. No arcs were introduced or forgotten so the set $W$ remains the same. The formal description of the recursive formula follows.
Given a reachability arrangement $(R,\ell)$ on $X_{t}$, let $(R^{\prime},\ell^{\prime})$ be the reachability arrangement induced by $X_{t^{\prime}}$ except that for each pair of vertices $u,w\in X_{t^{\prime}}$ where $(u,w)\notin E(R)$, $(u,v)\in E(R)$, and $(v,w)\in E(R)$ such that $\ell(u,v)=\ell(v,w)=\mathrm{future}$ we have $(u,w)\in E(R^{\prime})$ and $\ell^{\prime}(u,w)=\mathrm{future}$.
$A[t,R,\ell,b,W]=\begin{cases}\infty&\text{if there exists }u\in X_{t^{\prime}}\text{ such that}\\ &\qquad(u,v)\in E(R)\text{ and }\ell(u,v)\neq\textrm{future},\\ \infty&\text{if there exists }u\in X_{t^{\prime}}\text{ such that}\\ &\qquad(v,u)\in E(R)\text{ and }\ell(v,u)\neq\textrm{future},\\ \infty&\text{if }b(v)\neq 0,\\ A[t^{\prime},R^{\prime},\ell^{\prime},b|_{X_{t^{\prime}}},W]&\text{otherwise}.\end{cases}$
Clearly this entry can be calculated in polynomial time given the previous table entries.
We can formally prove the correctness of this formula by considering the family of sets compatible with $R,\ell,b,W$. In the first three cases this family is empty since $v$ is isolated in $G_{t}$. In the final case the families considered by the two entries are the same.
As the arc set in $G_{t}$ did not change, the parts 1, 2, 3, and 6 of Definition 3.15 remain true. Part 5 is the same for $X_{t}$ and we argued why $b(v)$ must be zero. For part 4, paths represented by future arcs in $R^{\prime}$ may stop satisfying the requirement to contain no vertices from $X_{t}\setminus\{u,w\}$; this is exactly the purpose of the modification in the formula.
Introduce arc node:
Assume $t$ introduces arc $(u,v)$ and its child is $t^{\prime}$. In any case, if $u$ and $v$ are in different strongly connected components then the new arc is inactive so it does not influence active degrees. We recognize two distinct cases based on whether this new arc belongs to $S$. On one hand, say the new arc $(u,v)\notin S$, then it may realize a future path from $u$ to $v$. Also, if $u$ and $v$ are in the same strongly connected component, then the added $(u,v)$ arc changes the active imbalance of $u$ and $v$ by one in $G_{t}$ but also in $G_{t}-S$ so the offset imbalance remains the same. On the other hand, if $(u,v)\in S$, then the active degree of its endpoints changes in $G_{t}$ but it does not change in $G_{t}-S$, hence, the offset imbalance changes by one. Note that the introduced arc $(u,v)$ may be one among multiple parallel copies of a multiarc – the only minor difference if we did not allow multiarcs would be to not allow the label on $(u,v)$ in $t^{\prime}$ to be $\mathrm{direct}$.
Let $\ell_{a}$ be the function such that $\ell_{a}(u,v)=a$ and $\ell_{a}(e)=\ell(e)$ for all other $e\in E(R)$. Let $b^{\prime}(u)=b(u)-1$, $b^{\prime}(v)=b(v)+1$, and $b^{\prime}(w)=b(w)$ for all $w\in X_{t^{\prime}}\setminus\{u,v\}$. Let $C_{u}$ be the strongly connected component of $R$ containing $u$. Let $W^{\prime}=W\setminus\{(u,v)\}$.
$A[t,R,\ell,b,W]=\begin{cases}A[t^{\prime},R,\ell,b,W^{\prime}]+1&\text{if }(u,v)\in W\text{ and }C_{u}\neq C_{v},\\ A[t^{\prime},R,\ell,b^{\prime},W^{\prime}]+1&\text{if }(u,v)\in W\text{ and }C_{u}=C_{v},\\ \min_{a}A[t^{\prime},R,\ell_{a},b,W]&\text{for all }a\in\{\mathrm{past},\mathrm{direct},\mathrm{future}\}\\ &\text{if }(u,v)\in E(R)\text{ and }\ell(u,v)=\mathrm{direct},\\ \infty&\text{otherwise}.\\ \end{cases}$
Again, this entry can be calculated in polynomial time given the previous table entries.
To prove correctness we consider a set $S$ compatible with $R,\ell,b,W$. When $(u,v)\in W$, clearly every set $S$ compatible with $R,\ell,b,W$ will contain $(u,v)$ by part 1 of Definition 3.15. Let $S^{\prime}=S\setminus\{(u,v)\}$, clearly this is compatible with $W^{\prime}$. Furthermore $G_{t}-S=G_{t^{\prime}}-S^{\prime}$ so $S^{\prime}$ satisfies parts 2, 3, 4, 6 with $R,\ell$. When $u$ and $v$ are in different components in $R$, $b_{G_{t}}^{R}(w)=b_{G_{t^{\prime}}}^{R}(w)$ for each $w\in X_{t}$ so part 5 is satisfied with the same $b$ and hence $S^{\prime}$ is compatible with $R,\ell,b,W^{\prime}$. Otherwise when $u$ and $v$ are in the same component in $R$, $b_{G_{t}}^{R}(u)=b_{G_{t^{\prime}}}^{R}(u)-1$ and $b_{G_{t}}^{R}(v)=b_{G_{t^{\prime}}}^{R}(v)+1$ so $S^{\prime}$ is compatible with $R,\ell,b^{\prime},W^{\prime}$. We move on to the case where $(u,v)\notin W$ but it is a direct arc in $(R,\ell)$. No compatible set $S$ considered here contains $(u,v)$ so the active imbalance of any vertex in $G_{t}$ and $G_{t}-S$ changes by the same amount and hence part 5 is still satisfied. Also $G_{t}-S=(G_{t^{\prime}}\cup\{(u,v)\})-S$. Since $(u,v)\in E(R)$, we have $(G_{t}-S)\cup R=(G_{t^{\prime}}-S)\cup R$ so part 6 is still satisfied. Parts 1, 2, 3, and 4 are clearly unchanged except on $(u,v)$ which is part of $G_{t}$ but not $G_{t^{\prime}}$ so we consider any set compatible with $R,\ell_{a},b,W$ for any $a\in\{\mathrm{past},\mathrm{direct},\mathrm{future}\}$.
Forget node:
If $t$ is a forget node with child $t^{\prime}$ such that $X_{t}=X_{t^{\prime}}\setminus\{v\}$ then we need to ensure that the forgotten vertex has zero active imbalance in $G_{t}-S$ and that there are no future arcs incident to it in the old arrangement. Zero active imbalance is equivalent to an offset imbalance of $-b_{G_{t}}^{R}(v)$, which we have precalculated. Also, the only change to the remaining reachability arrangement should be new past arcs where there was previously a path through $v$.
Let $\mathcal{R}$ be the set of reachability arrangements $(R^{\prime},\ell^{\prime})$ on $X_{t^{\prime}}$ such that
  1. 1.
    for every arc $e\in E(R^{\prime})$ incident to $v$ we have $\ell^{\prime}(e)\neq\mathrm{future}$,
  2. 2.
    if $(u,v),(v,w)\in E(R^{\prime})$ while either $(u,w)\notin E(R^{\prime})$ or $\ell^{\prime}(u,w)\neq\mathrm{direct}$, then $(u,w)\in E(R)$ and $\ell(u,w)=\mathrm{past}$,
  3. 3.
    for all pairs of vertices $u,v\in X_{t^{\prime}}$ that are not resolved in the previous points, $(u,w)\in E(R^{\prime})\Leftrightarrow(u,w)\in E(R)$ and $\ell^{\prime}(u,w)=\ell(u,w)$.
Let $\mathcal{Q}$ be a set of quadruplets $(R^{\prime},\ell^{\prime},b^{\prime},W^{\prime})$ be such that the reachability arrangement is picked out of the set described above $(R^{\prime},\ell^{\prime})\in\mathcal{R}$, the active imbalance is the same for all vertices but $v$ as described, i.e., $b^{\prime}|_{X_{t}}=b$ and $b^{\prime}(v)=-b_{G_{t}}^{R}(v)$, and part of the solution on $X_{t}$ is the same as before except for the forgotten vertex $W^{\prime}[X_{t}]=W$. Then
$A[t,R,\ell,b,W]=\min{\{A[t^{\prime},R^{\prime},\ell^{\prime},b^{\prime},W^{\prime}]:(R^{\prime},\ell^{\prime},b^{\prime},W^{\prime})\in\mathcal{Q}\}}.$
We have no more than $4^{{\mathtt{tw}}^{2}}$ reachability arrangements in $\mathcal{R}$ which can be easily iterated through brute-force, $b^{\prime}$ is uniquely determined by $b$ and $b_{G_{t}}^{R}(v)$ (which is precalculated), and there are no more than $2^{{\mathtt{tw}}}$ possibilities of how $W^{\prime}$ can look like. Hence, we can calculate $A[t,R,\ell,b,W]$ in time $2^{{\mathtt{tw}}}\cdot 4^{{\mathtt{tw}}^{2}}\cdot n^{{\mathcal{O}}(1)}$.
To prove correctness we consider an $S$ that is compatible with $R,\ell,b,W$, clearly $S$ is also compatible with some $W^{\prime}$ at $t^{\prime}$ (part 1 of Definition 3.15). Being compatible with $b^{\prime}$ ensures that $v$ has offset imbalance such that it has active imbalance zero and every other vertex in $V(G_{t^{\prime}})\setminus X_{t^{\prime}}$ is also in $V(G_{t})\setminus X_{t}$ so part 6 is satisfied. Part 5 is satisfied since all the other offset imbalances do not change. For any $(R^{\prime},\ell^{\prime})\in\mathcal{R}$, the first condition on $\mathcal{R}$ ensures that $G_{t}-S$ does not need any paths through the future to $v$ (part 4): necessary since $v$ is in the past in $G_{t}$. The second and third conditions ensure that the other changes to the reachability arrangement caused by $v$ not being part of $X_{t}$ are appropriate and hence parts 2 and 3 are satisfied.
Join node:
When merging two nodes $t_{1}$ and $t_{2}$ to a parent join node $t$ the reachability arrangements should be nearly the same. The notable exception is that past arcs in the parent arrangement can be either past in both child arrangements or we can have past arc in one arrangement while there is a future arc on the other arrangement. In a similar way, we need to consider for each $u\in X_{t}$ how the imbalance $b(u)$ in $G_{t}$ is made up of parts in $G_{t_{1}}$ and $G_{t_{2}}$. The new compatible solutions are unions of the solutions compatible with pairs of such arrangements. Their overlap is exactly $W$ so the size of the union is simply the sum of their sizes minus $|W|$.
Given a reachability arrangement $(R,\ell)$ let $L_{R,\ell}$ be the set of pairs of functions $\ell_{1},\ell_{2}$ such that for all $e\in E(R)$
  1. 1.
    if $\ell(e)=\mathrm{past}$, then $(\ell_{1}(e),\ell_{2}(e))\in\{(\mathrm{past},\mathrm{past}),(\mathrm{future},\mathrm{past}),(\mathrm{past},\mathrm{future})\}$,
  2. 2.
    otherwise $\ell_{1}(e)=\ell_{2}(e)=\ell(e)$.
Note that because arcs incident to $u$ can be partitioned to those having the other endpoint in $X_{t}$ and those with the other endpoint in $G_{t}-X_{t}$ we have $b_{G_{t}}^{G-S}(u)=b_{G_{t}[X_{t}]}^{G-S}(u)+b_{G_{t}-(X_{t}\setminus u)}^{G-S}(u)$, in the same way we can decompose $b_{G_{t}-S}^{G-S}(u)$. For each $u\in X_{t}$ we can unpack its imbalance to get the following.
$b(u)$ $={\mathrm{off}}_{G_{t}-S,G_{t}}^{R}(u)={\mathrm{off}}_{G_{t}-S,G_{t}}^{R}(u)=b_{G_{t}}^{R}(u)-b_{G_{t}-S}^{R}(u)$
$=\big{(}b_{G_{t}[X_{t}]}^{R}(u)+b_{G_{t}-(X_{t}\setminus u)}^{R}(u)\big{)}-\big{(}b_{(G_{t}-S)[X_{t}]}^{R}(u)+b_{(G_{t}-S)-(X_{t}\setminus u)}^{R}(u)\big{)}$
$=\big{(}b_{G_{t}[X_{t}]}^{R}(u)-b_{(G_{t}-S)[X_{t}]}^{R}(u)\big{)}+\big{(}b_{G_{t}-(X_{t}\setminus u)}^{R}(u)-b_{(G_{t}-S)-(X_{t}\setminus u)}^{R}(u)\big{)}$
$={\mathrm{off}}_{G_{t}[X_{t}]-S,G_{t}[X_{t}]}^{R}(u)+{\mathrm{off}}_{G_{t}-(X_{t}\setminus u)-S,G_{t}-(X_{t}\setminus u)}^{R}(u)$
Note that this decomposition works the same if we consider $G_{t_{1}}$ or $G_{t_{2}}$ instead of $G_{t}$. Importantly, we have $G_{t_{1}}[X_{t_{1}}]=G_{t_{2}}[X_{t_{2}}]$ so the first term is equivalent with respect to both child nodes $t_{1}$ and $t_{2}$. The second term counts active degree that is exclusive to each child, hence, let $b_{1}$ and $b_{2}$ be functions such that for each $u\in X_{t}$ we have
$b(u)=b_{1}(u)+b_{2}(u)-{\mathrm{off}}_{G_{t}[X_{t}]-S,G_{t}[X_{t}]}^{R}(u).$
The term ${\mathrm{off}}_{G_{t}[X_{t}]-S,G_{t}[X_{t}]}^{R}(u)$ is computable in polynomial time. Let $\mathcal{B}$ be the set of $(b_{1},b_{2})$ pairs that conform to the above equality.
The compatible arc sets overlap on $G_{t}[X_{t}]$ so we compute the entries as follows
[…]
$:(\ell_{1},\ell_{2})\in L_{R,\ell},(b_{1},b_{2})\in\mathcal{B}\}.$
$L_{R,\ell}$ contains at most $3^{{\mathtt{tw}}^{2}}$ pairs of functions. Functions in $\mathcal{B}$ are defined over range $[-\alpha,\alpha]$ and for a fixed $b$ there is at most $12\alpha+$ 1 ways to choose $b_{1}(u)$ which fixes $b_{2}(u)$. As we can choose the values of these functions for each $u\in X_{t}$ independently there is at most $(2\alpha+1)^{{\mathtt{tw}}}$ ways to choose suitable $b_{1}$ and $b_{2}$. The minimum is simply over $L_{R,\ell}$ and $\mathcal{B}$ so this entry can be calculated in $3^{{\mathtt{tw}}^{2}}(2\alpha+1)^{{\mathtt{tw}}}\cdot n^{{\mathcal{O}}(1)}$ time.
To prove the correctness of this formula we consider a set $S$ compatible with $R,\ell,b,W$. Let $S_{1}=S\cap E(G_{t_{1}})$, $S_{2}=S\cap E(G_{t_{2}})$. Clearly both $S_{i}$ satisfy part 1 of Definition 3.15 since their intersection with the bag is the same as $S$. Similarly part 2 is the same in $t$, $t_{1}$ and $t_{2}$ since nothing changes on the bag. Paths through the past in $G_{t}-S$ pass through exactly one of $G_{t_{1}}$ and $G_{t_{2}}$ so there must be a path through the past in at least one of the $G_{t_{i}}-S_{i}$s (there may be one in both since part 3 only requires that such a path exists). Paths through the future (part 4) in $G_{t}-S$ are also through the future in $G_{t_{i}}-S_{i}$ since $G_{t_{i}}\subseteq G_{t}$. The active imbalances are calculated with respect to $G_{t}-S\cup R$. Since $G_{t_{2}}-X_{t}$ is the future from the perspective of $t_{1}$ all the paths in $G_{t_{2}}-X_{t}-S_{2}$ are represented by arcs of $R$. Hence $(G_{t}-S)\cup R$ has the same strongly connected components on $X_{t}$ as $(G_{t_{i}}-S_{i})\cup R$ and therefore the active imbalance of each vertex in $G_{t_{i}}-X_{t}$ is zero by part 6. Finally the offset imbalance is shared; this is exactly the purpose of $\mathcal{B}$.
For a fixed node $t$ there are $4^{{\mathtt{tw}}^{2}}$ reachability arrangements on $X_{t}$, $(2\alpha+1)^{{\mathtt{tw}}}$ possible $b$’s, and $2^{{\mathtt{tw}}^{2}}$ possible $W$’s. Both introduce vertex and introduce arc node compute their entry from a fixed entry of their child node in $n^{{\mathcal{O}}(1)}$ time. Forget node is computed in $2^{{\mathtt{tw}}}\cdot 4^{{\mathtt{tw}}}\cdot n^{{\mathcal{O}}(1)}$ while join node is computed in $3^{{\mathtt{tw}}^{2}}\cdot(2\alpha+1)^{{\mathtt{tw}}}\cdot n^{{\mathcal{O}}(1)}$ time.
It is known that the total number of nodes in the nice tree decomposition with introduce arc nodes is $n^{{\mathcal{O}}(1)}$ and it can be observed that this still holds for the extension on multiarcs. Hence, the overall run time is
$\big{(}4^{{\mathtt{tw}}^{2}}\cdot(2\alpha+1)^{{\mathtt{tw}}}\cdot 2^{{\mathtt{tw}}^{2}}\big{)}\cdot\big{(}2^{{\mathtt{tw}}}\cdot 4^{{\mathtt{tw}}}+3^{{\mathtt{tw}}^{2}}\cdot(2\alpha+1)^{\mathtt{tw}}\big{)}\cdot n^{{\mathcal{O}}(1)}=24^{{\mathtt{tw}}^{2}}\cdot(2\alpha+1)^{2{\mathtt{tw}}}\cdot n^{{\mathcal{O}}(1)}$
4 Our Results for ESCAD on Simple Digraphs
In this section, we study ESCAD on simple digraphs, which we formally define as follows.
Simple Eulerian Strong Component Arc Deletion (SESCAD) Input: A simple digraph $G$, an integer $k$ Question: Is there a subset $R\subseteq E(G)$ of size $|R|\leq k$ such that in $G-R$ each strongly connected component is Eulerian?
Let us begin by stating a simple observation that enables us to make various inferences regarding the complexity of SESCAD based on the results we have proved for ESCAD.
Observation 4.1.
Consider an ESCAD instance $\mathcal{I}=(G,k)$. If we subdivide every arc $(u,v)$ into $(u,w),(w,v)$ (using a new vertex $w$) then we get an equivalent SESCAD instance $\mathcal{I}^{\prime}=(G^{\prime},k)$ with $|V(G^{\prime})|=|V(G)|+|E(G)|$ and $|E(G^{\prime})|=2|E(G)|$. Moreover, each arc of the solution to $\mathcal{I}$ is mapped to one respective arc of the subdivision and vice versa.
4.1 Hardness Results for SESCAD
We first discuss the implications of Theorems 1.1, 1.2 and 1.4 for SESCAD along with 4.1.
Corollary 4.2.
SESCAD is W$[1]$-hard when parameterized by the solution size.
Proof.
Follows from Theorem 1.1 and 4.1. ∎
Observation 4.3.
If we subdivide all arcs in a digraph $G$ that has a vertex cover $X$, we get a simple digraph $G^{\prime}$ such that $G^{\prime}-X$ is the disjoint union of directed stars.
Corollary 4.4.
SESCAD is W$[1]$-hard parameterized by minimum modulator size to disjoint union of directed stars.
Using the stronger assumption of ETH, we have the following result.
Theorem 4.5.
There is no algorithm solving SESCAD in $f(k)\cdot n^{o(k/\log k)}$ time for some function $f$, where $k$ is the size of the smallest vertex set that must be deleted from the input graph to obtain a disjoint union of directed stars and $n$ is the input length, unless the Exponential Time Hypothesis fails.
Proof.
The reduction in the proof of Theorem 1.2 along with Propositions 3.8, 4.1 and 4.3 implies the statement. ∎
Note that the above result rules out an FPT algorithm for SESCAD parameterized by various width measures such as treewidth and even treedepth.
Theorem 4.6.
[…]
Proof.
Follows from Theorem 1.4 and 4.1. ∎
4.2 FPT Algorithms for SESCAD
Firstly, the FPT algorithms discussed in the previous section naturally extend to SESCAD. However, for SESCAD, the lower bound parameterized by modulator to a disjoint union of directed stars leaves open the question of parameterizing by larger parameters. For instance, the vertex cover number.
To address this gap, we provide an FPT algorithm for SESCAD parameterized by vertex integrity, a parameter introduced by Barefoot et al. [1].
Definition 4.7 (Vertex Integrity).
An undirected graph $G=(V,E)$ has vertex integrity $k$ if there exists a set of vertices $M\subseteq V$, called a $k$-separator, of size at most $k$ such that when removed each connected component has size at most $k-|M|$. A directed graph has vertex integrity $k$ if and only if the underlying undirected graph has vertex integrity $k$. The notion of a $k$-separator in digraphs carries over naturally from the undirected setting.
FPT algorithms parameterized by vertex integrity have gained popularity in recent years due to the fact that several problems known to be W$[1]$-hard parameterized even by treedepth can be shown to be FPT when parameterized by the vertex integrity [9]. Since Corollary 4.4 rules out FPT algorithms for SESCAD parameterized by treedepth, it is natural to explore SESCAD parameterized by vertex integrity and our positive result thus adds SESCAD to the extensive list of problems displaying this behavior.
Moreover, this FPT algorithm parameterized by vertex integrity implies that SESCAD is also FPT when parameterized by the vertex cover number and shows that our reduction for ESCAD parameterized by the vertex cover number requires multiarcs for fundamental reasons and cannot be just adapted to simple digraphs with more work.
We will use as a subroutine the well-known FPT algorithm for ILP-Feasibility. The ILP-Feasibility problem is defined as follows. The input is a matrix $A\in{\mathbb{Z}}^{m\times p}$ and a vector $b\in{\mathbb{Z}}^{m\times 1}$ and the objective is to find a vector $\bar{x}\in{\mathbb{Z}}^{p\times 1}$ satisfying the $m$ inequalities given by $A$, that is, $A\cdot\bar{x}\leq b$, or decide that such a vector does not exist.
Proposition 4.8 ([13, 14, 8]).
ILP-Feasibility can be solved using ${\mathcal{O}}(k^{2.5k+o(k)}\cdot L)$ arithmetic operations and space polynomial in $L$, where $L$ is the number of bits in the input and $k$ is the number of variables.
See 1.6
Proof.
Consider an instance $(G,p)$ of SESCAD, where $G$ has vertex integrity at most $k$. Suppose that this is a yes-instance with a solution $S$ and let $M$ be a $k$-separator of $G$. Without loss of generality, assume that $V(G)=[n]$ and $M=[|M|]$. In our algorithm, we only require the fact that since $M$ is a $k$-separator in a digraph $G$, every weakly connected component of $G-M$ has size at most $k$ (recall, the definition of vertex integrity bounds the component sizes even more). Further, we remark that our algorithm does not require a $k$-separator to be given as input since there is an FPT algorithm parameterized by $k$ to compute it [7].
We next guess those arcs of $S$ that have both endpoints in $M$, remove them and adjust $p$ accordingly. The number of possible guesses is $2^{{\mathcal{O}}(k^{2})}$. Henceforth, we assume that every arc in the hypothetical solution $S$ has at least one endpoint disjoint from $M$.
We next guess the reachability relations between the vertices of $M$ in $G-S$. The correct guess is called the reachability signature of $M$ in $G-S$, denoted by $\sigma$, which is a set of ordered pairs where, for every $m_{1},m_{2}\in M$, $(m_{1},m_{2})\in\sigma$ if and only if $m_{2}$ is reachable from $m_{1}$ in $G-S$. The number of possibilities for $\sigma$ is clearly bounded by $2^{{\mathcal{O}}(k^{2})}$.
For every simple digraph comprised of at most $|M|+k$ vertices and every possible injective mapping $\lambda$ of $M$ to the vertices of this digraph, we define the type of this digraph as the label-preserving isomorphism class with the labeling $\lambda$. Denote the set of all types by $\mathsf{Types}$. For each type $\tau\in\mathsf{Types}$, we denote by $G_{\tau}$ a fixed graph of this type that we can compute in time depending only on $k$. Due to the labeling injectively mapping $M$ to the vertices of $G_{\tau}$, we may assume that $M\subseteq V(G_{\tau})$.
The number of types is clearly bounded by a function of $k$ and for each weakly connected component (from now onwards, simply called a component) $C$ of $G-M$ and graph $G_{C}=G[C\cup M]$ with the vertices of $M$ mapped to themselves by the identity labeling on $M$, denoted $\lambda_{M}$, we compute the type of the graph $G_{C}$. From now on, we drop the explicit reference to $\lambda_{M}$ as it will be implied whenever we are handling the graph $G_{C}$. For every type $\tau$, we also compute the number $n_{\tau}$ of components $C$ such that $G_{C}$ is of type $\tau$. Since the type of each $G_{C}$ can be computed in $f(k)$-time for some function $f$, this step takes FPT time.
Following that, for every component $C$, and every arc set $S_{C}$ in $G_{C}$, we check whether the type of $G_{C}-S_{C}$ (with labeling $\lambda_{M}$) is compatible with $\sigma$. To be precise, for a set $S_{C}$ of arcs in the digraph $G_{C}$, we verify that every vertex of $C$ is balanced in its strongly connected component in the graph $G^{\prime}_{C}=G_{C}-S_{C}+\sigma$. If the answer to this check is yes, then this is a compatible type. Notice that by adding the ordered pairs in $\sigma$ as arcs to $G_{C}-S_{C}$, we ensure that the arcs of the graph we take into account in this check on balances of the vertices in $C$ (i.e., active arcs) are exactly all those arcs that are already in $\mathrm{strong}(G_{C}-S_{C})$ plus those arcs of $G_{C}-S_{C}$ that would be inside a strongly connected component if the relations in $\sigma$ were realized. Since each component $C$ has size bounded by $k$, the number of possibilities for $S_{C}$ is bounded by a function of $k$ for each component (here, we crucially use the fact that we have a simple digraph), and hence, in FPT time, we can compute a table $\Gamma$ stating, for every $C$ and $S_{C}$ subset of arcs in $G_{C}$, whether the type of $G_{C}-S_{C}$ is compatible with $\sigma$.
Notice that for each component, deleting the arcs of the hypothetical solution $S$ from each component $C$ transitions $G_{C}$ from one type to another type that is compatible with $\sigma$. To be precise, for each $C$ and set $S_{C}=S\cap A(G_{C})$, we can think of $S_{C}$ as taking $G_{C}$ from the type of $G_{C}$ (call it $\tau_{1}$) to the type of $G_{C}-S_{C}$ (call it $\tau_{2}$), at cost $|S_{C}|$. Moreover, the type $\tau_{2}$ is compatible with $\sigma$. Thus, the table $\Gamma$ encodes the cost of transitioning each graph $G_{C}$ to a type compatible with $\sigma$. This can be expressed by a value $\mathrm{cost}(\tau_{1},\tau_{2})$ for every pair of types. If $\tau_{2}$ is not compatible with $\sigma$, then set this value to be prohibitively high, say the number of arcs in $G$ plus one. Otherwise, $\mathrm{cost}(\tau_{1},\tau_{2})$ is given by the table $\Gamma$.
In our next step, we guess a set of $O(k^{2})$ types such that for every pair of vertices $m_{1},m_{2}\in M$, if $\sigma$ requires that $m_{1}$ can reach $m_{2}$, then there is a sequence of vertices of $M$ starting at $m_{1}$ and ending in $m_{2}$ such that for every consecutive ordered pair $(x,y)$ in this sequence, either $(x,y)$ is an arc in $G[M]$ (and since it is not already deleted, it is disjoint from $S$) or there is an $x$-$y$ path with all internal vertices through a subgraph that belongs to one of these ${\mathcal{O}}(k^{2})$ types. Call this set of types $T^{*}$. The bound on the size of $T^{*}$ comes from the fact that there are ${\mathcal{O}}(k^{2})$ pairs in $\sigma$.
Finally, whether or not the vertices of $M$ are balanced in ${\mathrm{s}trong}(G-S)$ is determined entirely by the number of graphs of each type in $G-S$ subject to the types in $T^{*}$ occurring. So, for every type, we determine the imbalance imposed by the type on each vertex of $M$ (taking $\sigma$ into account). To be precise, for every type $\tau$ and vertex $u\in M$, the imbalance on $u$ due to $\tau$ is denoted by $I(\tau,u)$ and is obtained by subtracting the number of active incoming arcs on $u$ from the number of active outgoing arcs on $u$, where an arc $(p,q)\in A(G_{\tau})$ where $u\in{p,q}$ is active, if and only if it lies in the same strongly connected component as $u$ in the graph $G_{\tau}+\sigma$.
All of the above requirements can be formulated as an ILP-Feasibility instance with $f(k)$ variables that effectively minimizes the total costs of all the required type transitions. More precisely, for every pair of types $\tau_{1}$ and $\tau_{2}$, we have a variable $x_{\tau_{1},\tau_{2}}$ that is intended to express the number of graphs $G_{C}$ of type $\tau_{1}$ that transition to type $\tau_{2}$. We only need to consider variables $x_{\tau_{1},\tau_{2}}$ where $\tau_{1}$ is the type of some $G_{C}$ and $\tau_{2}$ is compatible with $\sigma$. So, we restrict our variable set to this. Moreover, for every $\tau$ that is compatible with $\sigma$, we have a variable $y_{\tau}$ that is intended to express the number of components $C$ such that $G_{C}$ transitions to type $\tau$.
Then, we have constraints that express the following:
  1. 1.
    The cost of all the type transitions is at most $p$.
    $\sum_{\tau_{1},\tau_{2}\in\mathsf{Types}}{\mathrm{c}ost}(\tau_{1},\tau_{2})\cdot x_{\tau_{1},\tau_{2}}\leq p$
  2. 2.
    For each type $\tau$ in $T^{*}$, there is at least one transition to $\tau$. This will ensure that the reachability relations required by $\sigma$ are achieved.
    $\sum_{\tau_{1}\in\mathsf{Types}}x_{\tau_{1},\tau}\geq 1$
  3. 3.
    For every component $C$, $G_{C}$ transitions to some type compatible with $\sigma$. So, for every type $\tau$, we have:
    $\sum_{\tau_{2}\in\mathsf{Types}}x_{\tau,\tau_{2}}=n_{\tau}$
    Recall that $n_{\tau}$ denotes the number of components $C$ such that $G_{C}$ is of type $\tau$ and we have computed it already.
  4. 4.
    The number of components $C$ such that $G_{C}$ transitions to type $\tau$, is given by summing up the values of $x_{\tau_{1},\tau}$ over all possible values of $\tau_{1}$.
    $\sum_{\tau_{1}\in\mathsf{Types}}x_{\tau_{1},\tau}=y_{\tau}$
  5. 5.
    The total imbalance imposed on each vertex of $M$ by the existing arcs incident to it, plus the imbalance imposed on it by the types to which we transition, adds up to 0.
    For each $u\in M$, let $\rho_{u}$ denote the imbalance on $u$ imposed by those arcs of $G[M]$ that are incident to $u$ and active in the graph $G[M]+\sigma$. The imbalance imposed on $u$ by a particular type $\tau$ is $I(\tau,u)$ and this needs to be multiplied by the number of “occurrences” of this type after removing the solution, i.e., the value of $y_{\tau}$.
    Hence, we have the following constraint for every $u\in M$.
    $\rho_{u}+\sum_{\tau\in\mathsf{Types}}I(\tau,u)y_{\tau}=0$
  6. 6.
    Finally, we need the variables to all get non-negative values. So, for every $\tau_{1},\tau_{2}\in\mathsf{Types}$, we add $x_{\tau_{1},\tau_{2}}\geq 0$ and for every $\tau\in\mathsf{Types}$, $y_{\tau}\geq 0$.
It is straightforward to convert the above constraints into the form of an instance of ILP-Feasibility. Since the number of variables is a function of $k$, Proposition 4.8 can be used to decide feasibility in FPT time. From a solution to the ILP-Feasibility instance, it is also straightforward to recover a solution to our instance by using the table $\Gamma$. ∎
5 Conclusions
We have resolved the open problem of Cechlárová and Schlotter [2] on the parameterized complexity of the Eulerian Strong Component Arc Deletion problem by showing that it is W$[1]$-hard and accompanied it with further hardness results parameterized by the vertex cover number and max-degree of the graph. On the positive side, we showed that though the problem is inherently difficult in general, certain combined parameterizations (such as treewidth plus either max-degree or solution size) offer a way to obtain FPT algorithms.
Our work points to several natural future directions of research on this problem.
  1. 1.
    Design of (FPT) approximation algorithms for ESCAD?
  2. 2.
    ESCAD parameterized by the solution size is FPT on tournaments [3]. For which other graph classes is the problem FPT by the same parameter?
  3. 3.
    Our FPT algorithm for SESCAD parameterized by vertex integrity is only aimed at being a characterization result and we have not attempted to optimize the parameter dependence. So, a natural follow up question is to obtain an algorithm that is as close to optimal as possible.
  4. 4.
    For which parameterizations upper bounding the solution size is ESCAD FPT? For instance, one could consider the size of the minimum directed feedback arc set of the input digraph as a parameter. Notice that in the reduction of Theorem 1.1, we obtain instances with unboundedly large minimum directed feedback arc sets due to the imbalance gadgets starting at the vertex $s_{j}$ for some color class $j$ and ending at the vertices in $\{x_{u}\mid u\in{\rm color~{}class~{}}j\}$.